### v1

# Test-Time Training Scaling for Chemical Exploration in Drug Design

Abstract
Chemical language models for molecular design have the potential to find solutions to multi-parameter optimization problems in drug discovery via reinforcement learning (RL). A key requirement to achieve this is the capacity to ”search” chemical space to identify all molecules of interest. Here, we propose a challenging new benchmark to discover dissimilar molecules that possess similar bioactivity, a common scenario in drug discovery, but a hard problem to optimize. We show that a population of RL agents can solve the benchmark, while a single agent cannot. We also find that cooperative strategies are not significantly better than independent agents. Moreover, the performance on the benchmark scales log-linearly with the number of independent agents, showing a test-time training scaling law for chemical language models.
Machine Learning, ICML
[…]
1 Introduction
Identifying better, safer, and more efficacious drug candidates faster can save hundreds of millions of dollars (Bender & Cortés-Ciriano, 2021), and bring medicines to patients years quicker. However, the vastness of chemical space (Polishchuk et al., 2013) is a central challenge. It is intractable to computationally enumerate all drug-like molecules, let alone evaluate them, synthesize them and experimentally test them in the lab. Traditionally, virtual screening is conducted of virtual libraries that contain just tiny fractions of drug-like chemistry that can be purchased readily and experimentally tested. On the other hand, goal-directed generative AI can generate novel chemical structures with optimal predicted properties while implicitly searching much larger chemical spaces (Arús-Pous et al., 2019a). Thus, harnessing the full potential of generative AI to facilitate drug design will have considerable positive impacts on drug discovery as a whole.
To fully exploit generative AI potential, achieving a thorough exploration of chemical space is critical. The more diverse solutions identified the more advantageous it is. Starting a drug discovery campaign from multiple diverse hits can offer backup series with alternative chemical and biological profiles, decreasing risk by increasing shots at goal (Brown & Boström, 2018). Additionally, probing novel areas of chemical space can avoid already patented compounds (Shimizu et al., 2023). Furthermore, in practice, many generative models undergo ”mode collapse” (a narrowing of chemical space) when optimizing for particular predicted properties, particularly those that are predicted by machine learning models with limited applicability domains arising from epistemic uncertainty (Preuer et al., 2018; Renz et al., 2019).
Reinforcement learning (RL) is a family of optimization algorithms that can train an agent (a decision-making entity) to maximize a reward within a Markov decision process (Sutton & Barto, 2018). In generative molecular design, the generative model acts as the RL agent that learns to adapt molecule generation to maximize a given reward. One or more predicted molecular properties can be combined into an ’oracle’ that provides the reward. A range of generative models as RL agents for molecular design have been demonstrated including heuristic rules combined with deep Q-networks (Zhou et al., 2019), deep language model policy functions (Olivecrona et al., 2017; Ståhl et al., 2019), or deep graph network policy functions (Atance et al., 2022). However, it is the deep language model policy functions, also referred to as chemical language models (CLMs) (Segler et al., 2018; Grisoni, 2023; Skinnider et al., 2021), that have emerged as the most popular generative drug design algorithms (Olivecrona et al., 2017; Thomas et al., 2022b; Guo & Schwaller, 2024), as they have demonstrated top performance on multiple benchmarks (Brown et al., 2019; Polykovskiy et al., 2020; Thomas et al., 2022a).
CLMs are autoregressive language models such as recurrent neural networks (RNNs) or GPT-style models trained on next token prediction typically using 1D SMILES (Weininger, 1988) strings that represent the 2D graph of a molecule. Seminal works (Olivecrona et al., 2017; Popova et al., 2018) applied RL to CLMs resulting in an agent capable of generating de novo molecules with desirable properties. The body of work now surrounding these models includes algorithmic variations to improve the learning efficiency (Bjerrum et al., 2023; Thomas et al., 2022b; Guo & Schwaller, 2024; Thomas et al., 2025), the practical utility (Fialková et al., 2021; Thomas et al., 2024a), as well as experimental validation of workflows utilizing such models for de novo drug design (Korshunova et al., 2022; Quinn et al., 2024; Thomas et al., 2024b). Interestingly, REINFORCE (Sutton & Barto, 2018) based RL algorithms remain superior for CLMs (Bou et al., 2024), likely due to strong policy initialization from pre-training (Ahmadian et al., 2024).
In the field of Large Language Models (LLMS), Test-Time Training (TTT) has emerged as a successful approach to improve abstract reasoning (Akyürek et al., 2024). TTT is a strategy that updates the parameters of a pre-trained LLM on a specific test-time task, where they are reset to the pre-trained parameters for the next task. Hence, the focus is on the predictions of the temporarily updated LLM instead of the updated LLM with its new parameters. Moreover, scaling inference-time has shown to improve performance on benchmarks log-linearly with the number of samples generated at inference (Brown et al., 2024). Together, they have led to landmark improvements in abstract reasoning benchmarks such as the ARC Challenge (Chollet, 2019).
Multi-agent reinforcement learning (MARL) involves the cooperation, competition, or evolution of multiple agents acting in an environment to enhance learning (Buşoniu et al., 2010). This is a proven effective strategy to reach human or greater level performance in StarCraft (Vinyals et al., 2019) and capture the flag (Jaderberg et al., 2019), or to learn cooperative microeconomic strategies (Johanson et al., 2022). Most relevant to the work presented here is the capacity of a population of agents to increase the diversity of solutions via cooperative learning by enforcing divergent behavior between agents (Parker-Holder et al., 2020). Although MARL is effective in fields such as game theory, few examples exist applying MARL to generative molecular design.
In this work, we provide two central contributions:
  1. 1.
    We introduce a new hard de novo drug design benchmark focused on efficient Molecule Exploration (MolExp). Unlike existing benchmarks, MolExp emphasizes discovering structurally diverse molecules that retain similar bioactivity, acting as an exploration challenge where multiple distinct regions of chemical space get high rewards and all should be found.
  2. 2.
    We show that a population of independent agents is better at chemical space exploration on the benchmark than a single agent trained for a long time, even with exploration bonuses. We also show that performance on the benchmark scales with the number of agents in the population. Lastly, we investigate several cooperative strategies between multiple agents to try to enhance exploration efficiency.
2 Related Work
Balancing the exploration-exploitation trade-off is a constant challenge in RL, including RL for drug design (Langevin et al., 2024). Exploitation initially maximizes the reward quicker but may lead to a lower overall reward, while exploration may lead to higher overall reward and increased diversity of solutions but takes longer. Diversity has consistently been an important topic in generative molecular design and is increasingly included in the evaluation of performance (Thomas et al., 2022a; Renz et al., 2024). Several approaches have been proposed to increase diversity, for example, using dual networks to balance exploitation-exploration during sampling (Liu et al., 2019), penalizing the reward of molecules similar to previously generated ones (Blaschke et al., 2020), or increasing the reward of molecules that are dissimilar to those previously generated (Thiede et al., 2022). Recent work shows that random network distillation (RND) (Burda et al., 2018) is particularly effective at increasing molecular diversity with low overhead due to not requiring an explicit memory of previously generated molecules (Svensson et al., 2024; Thomas et al., 2025).
Previous work has applied multi-agent RL to de novo drug design. Hu et al. (2024) deployed up to 4 GPT agents as CLMs that use the REINVENT loss formulation with an additional penalty to discourage the $k$th agent from being similar to previous agents. This multi-agent setup outperformed baseline algorithms on the GuacaMol benchmark (Brown et al., 2019). However, this benchmark consists of mostly simple tasks (as also stated by the authors). As demonstrated by their experiment comparing 1, 2, 4 and 8 agents all having near maximum performance and all within error. Moreover, independent agents and other cooperative strategies were not investigated. Therefore it remains to be seen whether MARL improves performance significantly, by how much, and which cooperative strategies work best.
Generative molecular design benchmarks must reflect realistic scenarios in drug design for practical application, as gold standard evaluation via synthesis and experimental validation of molecules is too costly at scale. The GuacaMol benchmark (Brown et al., 2019) proposed a suite of 20 tasks, the authors highlighted that $\sim 15/20$ tasks are easily solved by the generative algorithms. The MolOpt benchmark limited these tasks to a budget of 10,000 molecules during optimization, shifting the focus to efficiency, however, this likely over-rewards exploitative algorithms as opposed to explorative ones. New criteria have been defined to additionally measure the quality of chemistry generated which has changed rank performance (Thomas et al., 2022a; Renz et al., 2024), moreover, rank performance on this MolOpt benchmark didn’t reflect performance on more realistic optimization objectives such as predicted binding affinity via docking (Bou et al., 2024). Hence, there remains significant scope for new exploration-focused objectives that better reflect realistic optimization objectives in drug design.
3 Methodology
3.1 Problem definition
The generative molecular design problem is the generation of molecules that are optimal in one or more properties defined by the drug discovery context, for example, drug-likeness and predicting binding affinity against a protein implicated in a disease. The scoring function (also referred to as an oracle) $s(x)\in[0,1]$ may combine one or more property predictors $p_{i}(x)$ that score a given molecule $x$ returning a real value $p_{i}:X\rightarrow\mathbb{R}$. Multi-parameter optimization can be achieved by combining multiple predicted properties, the simplest example being to take the arithmetic mean $\sum_{i}t_{i}(p_{i}(x))$ ensuring that $t(p_{i}(x))\in[0,1]$ by using a transformation function $t_{i}$. Overall, the final score reflects the desirability of the molecule for the given context, however, optimizing a score combining many properties is usually non-trivial as improving one property can be detrimental to another resulting in a Pareto-optimization problem.
$\max_{x\in X}s(x)$ (1)
[…]
$\exists x_{i},...,x_{N}\in X,\,x_{i}\neq x_{j}\text{ such that }s(x_{i})\geq T$ (2)
We propose that a generative model benchmark should adequately mimic this concept of identifying multiple, seemingly unrelated molecular targets in chemical space (Figure 1, right). Firstly, previously published benchmarks typically don’t capture this with objectives focused on a single narrow target, or a single broad target (Figure 1, left). These do not adequately test the ability of an algorithm to explore chemical space if it is already receiving high rewards. Secondly, there are post-hoc metrics that measure the diversity of solutions (Thomas et al., 2022a; Renz et al., 2024), however, this only measures the diversity inherently embedded in the objective. Therefore, designing an objective to contain multiple distinct targets provides a greater opportunity to observe inherent exploration capability. Lastly, previous benchmarks predominantly measure performance on the maximum $S(x)$ achieved, whereas, we propose to measure performance on maximum $p_{i}(x)$ for each molecular target. This focuses evaluation on the algorithm’s ability to explore all high-rewarding regions.
Figure 1: Schematic representation of chemical space manifold according to molecular desirability, where energy wells are desirable sub-spaces of ’high-reward’ areas according to a given objective. (Left) Typical objectives in the literature. (Right) Our proposal objective. The sphere represents a generative model learning to optimize desirability or maximize reward. Target rediscovery where the reward reflects maximal similarity to target molecule in chemical space. Target similarity where the reward reflects similarity above a certain threshold to a target molecule in chemical space, perhaps with additional multi-parameter objectives. Predicted property of molecules in chemical space, where multiple regions of reward exist. Lastly, our proposal, multiple target rediscovery where there exists different and distinct regions of high reward chemical space, and the goal is to rediscover all target molecules.
3.2 Reinforcement learning with chemical language models
CLMs utilize a language representation, commonly SMILES (Weininger, 1988), to leverage natural language models such as recurrent neural networks (RNNs) or GPT-style models. For example, SMILES represent the molecular structure of paracetamol with the string ”CC(=O)Nc1ccc(O)cc1”. These models operate autoregressively, predicting the next token $x_{t}$ in a sequence $X=(x_{1},x_{2},\ldots,x_{T})$ based on the context provided by previous tokens. The process can be formalized as maximizing the conditional probability of each token given the preceding sequence $P(X)=\prod_{t=1}^{T}P(x_{t}|x_{<t})$, where $x_{<t}$ denotes the tokens up to $t-1$. The model is trained using the cross-entropy loss function which is equivalent to the negative log-likelihood Equation 3. Where $T_{i}$ is the length of the $i$-th sequence, and $\theta$ are the model parameters.
$\mathcal{L}=-\sum_{t=1}^{T_{i}}\log P_{\theta}(x_{t}|x_{<t})$ (3)
Reinforcement learning (RL) offers a structured approach to molecular generation by formalizing the problem as a Markov Decision Process (MDP) defined by $\langle S,A,R,P,\rho_{0}\rangle$. Here, $S$ represents the states (i.e., partial molecular structures), $A$ denotes available actions (i.e., adding the next token that could describe an atom), $R$ defines a scalar reward function (i.e., evaluating molecular properties), $P$ is the state transition probability, and $\rho_{0}$ is the initial state distribution. Here, the pre-trained CLM acts as an already parameterized policy $\pi_{\theta}$ that defines an initial transition probability of selecting the next action at a given timestep $P(s_{t+1}|a_{t},s_{t})$.
The specific RL algorithm used is a REINFORCE-based policy-gradient algorithm. REINFORCE learns a policy $\pi_{\theta}$ that maximizes the expected cumulative reward $J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)]$, where $\tau$ represents a trajectory of states and actions. Its core objective is to adjust the policy parameters $\theta$ such that the likelihood of actions leading to higher rewards increases. This is achieved by updating $\theta$ proportionally to the gradient of the reward-weighted log-probability of action sequences, expressed as $\nabla J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\cdot R(\tau)\right]$. REINFORCE is particularly suited for problems with sparse or delayed rewards, as it treats entire trajectories as units of optimization. This works well when building molecules which only receive a reward when the building is complete.
We deviate to vanilla REINFORCE by reshaping the reward to regularize the agent policy to the prior policy $\pi_{prior}$ with a coefficient $\sigma$ and by adding an exponent $\alpha$, as proposed by Thomas et al. (2025). These components improve the quality of molecules proposed by the agent and improve performance by steepening the gradients in the reshaped reward landscape. In addition to reshaping the reward, we implement experience replay to augment on-policy data with off-policy data at each iteration. A replay buffer is kept of high-reward molecules which is sampled with prioritization proportional to their reward. This additionally improves learning efficiency (Gao et al., 2022; Thomas et al., 2025).
$R(\tau)_{reshaped}=clip(R(\tau)+\sigma\cdot\pi_{prior})^{\alpha}$ (4)
3.3 Multi-agent reinforcement learning
MARL extends the framework of single-agent RL to systems with multiple independent or interacting agents, each learning to optimize its own policy in an environment. The problem is the same as defined before, but now with $\mathcal{N}$ agents.
In the independent case, each agent still aims to maximize its cumulative reward using its own $\langle\mathcal{S}^{i},\mathcal{A}^{i},R^{i},P^{i}\rangle$ for $i\in\mathcal{N}$ with an independent replay buffer for each agent. Note that independent agents will still deviate due to the stochastic nature of the transition probabilities $P(s^{i}_{t+1}|a^{i}_{t},s^{i}_{t})$ leading to the sampling of different actions $a^{i}_{t}$.
In the interaction case, we focused on methods of cooperation within a population-learning setting which typically results in more diverse solutions (Oroojlooy & Hajinezhad, 2023). In this case, using $\langle\mathcal{S},\mathcal{A}^{i},R,P^{i}\rangle$ for $i\in\mathcal{N}$ where the states $S$ and $R$ may be shared across agents in a cooperative manner. For simplicity, here we refer to the complete molecule and final state $s_{T}$ simply as the state. We investigated the following methods of cooperation (for additional detail and equations see Appendix B):
Purge
Ensure agent replay buffers are mutually exclusive with respect to states contained within them.
Shared
Share a single replay buffer for all agents. This should result in policy convergence and act as a type of negative control.
Shared with bonus
Same as Shared but with a reward bonus for novel states not contained within the shared replay buffer.
Noise
Apply Gaussian noise to each agent’s policy parameters at the beginning of training.
RND
Apply RND universally to provide a reward bonus for novel states.
ENTS
Minimize the entropy of policy uncertainty across population states to encourage specialized behavior.
CES
Minimize the cross-entropy loss of policy uncertainty across population states to encourage explicit specialization towards its own states.
DIFFS
Maximize the difference between policy uncertainty in the agents own states relative to population states.
DIFFN
Re-implementation of the cooperative strategy applied by Hu et al. (2024) which seeks to maximize the log-probability difference of on-policy states between the current agent and other agents.
DvD
Diversity via Determinant (DvD) inspired by Parker Holder et al. (2020). Calculate a policy behavior embedding and maximize the divergence in policy behavior as measured by the determinant of the behavioral embedding transformed by a kernel function.
POPNORM
Normalize the agent return by subtracting the average return of the population for on-policy states, encouraging divergent behavior.
4 Experiments
To assess the ability of single and multiple RL agents to explore chemical space, we conducted three experiments.
  1. 1.
    We defined a molecule exploration benchmark and evaluated baseline algorithm performance.
  2. 2.
    We tested the effectiveness of different population sizes of multiple independent agents, and compared them to equivalent single agents.
  3. 3.
    We investigated different strategies of cooperation within a population of agents to enhance exploration efficiency.
4.1 Molecule exploration benchmark
To mimic our proposed problem formulation of identifying all hits in chemical space requiring intrinsic exploration, we designed a new molecular exploration (MolExp) benchmark. To embed this benchmark in real-world discovery, it was designed based on the converse of the similarity principle  [^1] In drug discovery, the similarity principle states that similar molecules $Sim(m_{1},m_{2})\geq\epsilon$ possess similar bioactivity $|f(m_{1})-f(m_{2})|\leq\delta$. However, this does not prove the converse and many exceptions exist that are more difficult to model., i.e., to identify multiple dissimilar molecules that exhibit similar desirable properties such as bioactivity. Note that activity cliffs are another example converse to the similarity principle that have been used to demonstrate more challenging scenarios for predictive models (Kwapien et al., 2022; Van Tilborg et al., 2022).
Each run is replicated on the benchmark 5 times and as with previous benchmarks (Gao et al., 2022) a default budget of 10,000 molecules is used.
Oracles The benchmark consists of four tasks, each using an oracle to evaluate molecules. These are based on drug candidates and their biological targets, including anti-psychotic drugs (AP), Adenosine A2A (A2A) receptor ligands, beta-secretase 1 (BACE1) inhibitors, and epidermal growth factor inhibitors (EGFR). These tasks contain 2, 3, 3, & 4 objectives respectively, where the objective is the similarity to a target molecule $sim(m_{i},m_{t})$. For each task, only a single reward is provided based on the similarity to the nearest target molecule $R(m)=max(sim(m_{i},m_{t_{1}}),...,sim(m_{i},m_{t_{N}}))$. Due to CLMs acting in string space, we chose a string-edit similarity function. This, reflects chemical space search scenarios where distinct regions of reward signal exist. A more detailed description of the molecular targets chosen and the similarity function can be found in Appendix C.
Metric Performance is measured by calculating the product of the top N most similar molecules to each target molecule respectively $\prod_{t\in T}\frac{\sum_{i\in N}sim(m_{i},m_{t})}{N}$. This metric focuses on the ability generate similar molecules to all objectives. Unless stated otherwise, we use $N=1$.
Baselines We tested virtual screening and multiple RL algorithms (see Appendix D) on their inherent ability for chemical exploration on the MolExp benchmark. Table 1 shows that the recently published ACEGENMolOpt (Thomas et al., 2025), performed best. Figure 13 illustrates that none of the tested algorithms optimize all molecular targets within the default budget (for a single replicate). Moreover, that the best scores are due to improved efficiency in maximizing similarity to a single target molecule. Therefore, there is still considerable room for improvement.
Estimated max An estimated maximum performance can be calculated for the chosen agent RL configuration (ACEGENMolOpt) by training the agent to maximize the similarity to each individual molecular target with the reward $R(m)=sim(m_{i},m_{t})$ only, Figure 14. This also tests the agents ability to optimize for the similarity to each individual molecular target in the first place which is an expected prerequisite for performance on the benchmark. Averaging the Top-1 similarity achieved over the 5 replicates and calculating the product over the group of molecular target objectives in each task, and then summing over all tasks led to an estimated achievable maximum benchmark score of 2.91.
Table 1: Baseline performance of virtual screening and policy-gradient based RL algorithms on the MolExp benchmark within a budget of 10,000 molecules. ACEGENMolOpt outperforms all other methods, while virtual screening outperforms two baselines algorithms vanilla REINFORCE and REINVENT.
Task SCREEN REINFORCE REINVENT REINVENTMolOpt AHC ACEGENPractical ACEGENMolOpt
AP 0.53 ± 0.04 0.56 ± 0.05 0.57 ± 0.04 0.63 ± 0.02 0.62 ± 0.06 0.64 ± 0.02 0.64 ± 0.04
A2A 0.38 ± 0.03 0.36 ± 0.02 0.39 ± 0.08 0.38 ± 0.03 0.40 ± 0.02 0.40 ± 0.07 0.44 ± 0.06
BACE1 0.29 ± 0.08 0.23 ± 0.02 0.21 ± 0.01 0.27 ± 0.03 0.23 ± 0.01 0.23 ± 0.03 0.28 ± 0.03
EGFR 0.17 ± 0.04 0.15 ± 0.02 0.16 ± 0.02 0.22 ± 0.03 0.17 ± 0.03 0.17 ± 0.02 0.25 ± 0.05
Sum 1.36 ± 0.10 1.30 ± 0.06 1.33 ± 0.09 1.50 ± 0.06 1.41 ± 0.07 1.44 ± 0.08 1.62 ± 0.09
4.2 Test-time training with RL for chemical exploration
We trained a population of independent agents using the best-performing baseline (ACEGENMolOpt) configuration. In this case, each agent had a budget of 10,000 molecules. The results of different population sizes up to 128 agents is shown in Figure 2. We found that the benchmark performance continued to increase log-linearly from $\sim 1.6$ to $\sim 3.5$ up until the maximum population size was tested, almost reaching the maximum of score of 4. Interestingly, when the population size reached 32, the estimated maximum score for this RL configuration was exceeded. This is indicative of the power of leveraging the variance between different agents to overcome the more ambiguous reward signal, which did not turn out to be a limiting factor. Figure 15 shows the average similarity to each target during training across multiple agents, indicating the variance observed between multiple agents with different agents maximizing similarity to different molecular targets, resulting in overall increased exploration and increased benchmark scores. This is exemplified in Figure 3 which shows that a population of 128 agents identifies all 3 A2A targets. Interestingly, for all population sizes in the AP task, all agents maximized the same molecular target, suggesting a particular difficulty in optimizing the other.
Figure 2: Benchmark performance on MolExp as the number of independent agents increases. The score reported is the product of the maximum similarity to each target molecule summed over the tasks. Note that performance increase scales log-linearly with increasing independent agents. Estimated max is calculated based on a single agent optimizing each task’s molecular target individually.
Figure 3: Example molecules generated during the A2A task from the MolExp benchmark. In columns, the molecule with the maximum similarity to each target is shown, alongside which agent sampled it. 1 agent almost solves target 3, 4 agents solve target 3 and almost target 2, meanwhile 128 agents solve all 3 targets.
Increasing the number of agents, each with an individual budget, effectively increases the total budget. Therefore, we tested the single agent with the cumulative budget of a population of agents, for comparison. Figure 4 shows that a single agent trained for the equivalent budget does not increase performance similarly. In this case, scaling the duration of training for a single agent results in a saturation of benchmark performance after approximately 40,000 molecules evaluated, as agents expectedly fail to continue exploring once a target has been maximized (Figure 16). This is only improved slightly if state-of-the-art exploration strategies are employed such as RND, the most performant exploration strategy (Svensson et al., 2024; Thomas et al., 2025), which provides an exploration bonus to the reward. Where Figure 17 shows only small fluctuations in similarity due to RND, before subsequent maximization again to the same ’chosen’ molecular target. This is instead of maximization of similarity to a given molecular target and then a switch from one to another, as intended with exploration methods.
Figure 4: Performance comparison between using multiple independent agents or running one single agent for longer. Note that running multiple independent agents with a budget of 10,000 each outperforms running a single agent with a combined budget of $10,000\times N_{agents}$. Additional exploration strategies such as RND don’t improve performance to the level of multiple independent agents.
4.3 Cooperative multi-agent RL for chemical exploration
Cooperation within a population of RL agents can theoretically improve performance by leveraging useful information shared between agents, such as, states already visited and their corresponding reward (Oroojlooy & Hajinezhad, 2023). We explored several strategies to leverage cooperative learning to improve the efficiency of chemical exploration. Considering the maximum number of task objectives in the benchmark was 4, we investigated strategies utilizing 4 RL agents. Where, the best-case scenario is if each agent learns a different objective.
The MolExp score of the different cooperative strategies are shown in Figure 5. All of the strategies tested were similar or worse in performance, with no statistically significant improvement over independent agents. The best strategies were POPNORM, ENTS, and DIFFN. Where lower $\lambda$ values performed best for ENTS showing divergent behavior for a single replicate in Figure 18 with little restriction in learning capability, although this is observed at larger values of $\lambda$. Meanwhile either the lowest (0.001) or highest (1) tested $\lambda$ value for DIFFN performed well. Interestingly, there didn’t seem to be any obvious correlation between the $\lambda$ value and restriction on learning capability for the values tested (Figure 20). Note that the best values here differ from the default of 0.01 used by Hu et al. (2024), also used with 4 CLM agents.
The Noise strategy exhibited higher standard deviation in benchmark score. Both Purge and Shared replay buffer strategies dropped performance, with Shared being among the worst performing benchmarks as anticipated. Where in the example replicate shown in Figure 18 every agent optimizes the same molecular target and hence converged in chemical space as anticipated, with the novelty bonus failing to ameliorate this effect. Purge may not offer strict enough segregation of replay buffers to diverge agent behavior, as unique but similar molecules may still exist in different replay buffers. On the other hand, strategies like RND and DvD that should more strongly encourage divergent behavior appear to ultimately restrict learning within the defined budget (Figure 18), without resulting in the desired outcome of different agents optimizing different molecular targets more than independent agents. Lastly, no tested values of $\lambda$ for CES or DIFFS outperformed independent agents. Where higher values also led to restricted learning capability, especially in the case of DIFFS as shown in Figure 19 and Figure 20.
Figure 5: Performance comparison of different multi-agent cooperative strategies on the MolExp benchmark. Each plot is a different cooperative strategy using 4 agents. The dashed line represents the average performance of 4 independent agents. Note that no strategy tested significantly outperformed a population of 4 independent agents as measured by a Bonferroni-corrected, one-tailed t-test.
5 Conclusion
In this paper, we present the challenge of exploring chemical space with a CLM to identify all possible candidate molecules in the context of drug discovery. We introduced a new benchmark to measure chemical exploration highlighting the poor intrinsic exploration of current baseline algorithms. We showed that this benchmark could be practically solved by a population size of up to 128 independent agents which performed superior to a single agent trained for longer, even with exploration bonuses. Lastly, the cooperative strategies we proposed and tested were unable to significantly outperform independent agents in our setting.
The scaling of populations of agents in our setting has a direct relation to the scaling of TTT in LLMs. In our case, a CLM agent is pre-trained to form a generalist model that can sample molecules in a similar chemical space to the training dataset. We temporarily update the CLM agent with RL on a specific test-time task (each of the four benchmark tasks), which can be viewed as iterations of training and sampling, and collect the outcomes (molecules sampled) but throw away the updated parameters. However, as opposed to scaling just the inference-time as typically done, we scaled the whole TTT process by increasing the number of agents undergoing TTT on a specific test-time task. Interestingly, observing the same log-linear scaling law in performance with the number of samples.
Overall, this work takes an important step towards the future scenario of identifying all possible molecules of interest in drug-like chemical space. Our MolExp benchmark is the first to explicitly ask this question, which we hope will guide algorithmic improvement until chemical space can be efficiently and thoroughly explored for a given context. Populations of RL agents are a promising solution towards this future.
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning in application to drug discovery. We don’t anticipate any direct societal consequences of our work that must be specifically highlighted here.
Appendix A Chemical language model implementation
We trained a gated recurrent unit (GRU) network on molecules within the ChEMBL34 database (Gaulton et al., 2012) to act as the CLM.
First, we filtered the database to form the training dataset and ensured no molecules existed with more 10 rotatable bonds, a logP above 5.5, a molecular weight outside of 150-650 Da, or contained atoms not in a comment set $a\in{C,N,S,O,F,Cl,Br,H}$. All molecules were neutralized if possible, converted to non-isomeric SMILES (i.e., containing no stereoinformation) and de-duplicated. This resulted in a dataset of 1,711,022 unique molecules. Finally, the target molecules of the MolExp benchmark Appendix C were purposefully confirmed to be present within the training dataset, to ensure that objectives were theoretically achievable.
The GRU consisted of an embedding of 256 and 3 layers of GRU cells with hidden dimension 512. This was trained with restricted SMILES randomization (Arús-Pous et al., 2019b) for 10 epochs with a batch size of 128. This was implemented such that a random SMILES string was each time a dataset molecule was sampled during training. An ADAM optimizer was employed with a learning rate of 0.001 and step scheduler that dropped the learning rate every 500 steps. The learning rate, loss and SMILES validity during training is shown in Figure 6.
Figure 6: Pre-training of the CLM on a training dataset of ChEMBL34 molecules.
Appendix B Cooperative strategies
We employed the following cooperative strategies using $\langle\mathcal{S},\mathcal{A}^{i},R,P^{i}\rangle$ for $i\in\mathcal{N}$ where the states $S$ and $R$ may be shared across agents in a cooperative manner. For brevity, we denote the probability of a full sequence generated as $\pi_{i}(\tau)$.
Purge
Purging each agent’s replay buffer to ensure no molecule exists in more than one replay buffer, resulting in a unique replay buffer for each agent.
Shared
One shared replay buffer was used by all agents as a method of communicating high-rewarding areas of chemical space by on-policy augmentation. We anticipate this to result in convergence of agents to the same areas of chemical space and hence act more as a negative control.
Shared with bonus
The same as Shared, however, an intrinsic reward $R(\tau)_{Nov}$ was applied as a novelty bonus for molecules not already contained within the replay buffer. This reward shaping was applied before Equation 4.
$R(\tau)_{reshaped}=\frac{R(\tau)+R(\tau)_{Nov}}{2}$ (5)
Noise
Gaussian noise was applied to each agent’s parameters once at initialization. Such as to perturb each agent’s starting policy to encourage learning divergence. The effect of $\lambda$ was tested on the validity of generated SMILES strings Figure 12. In cooperative experiments, we applied $\lambda_{Noise}=0.01$ noise to all GRU hidden layers and the final linear layer.
$\theta^{\prime}_{step=0}=\theta_{step=0}+\lambda\cdot\mathcal{N}(0,1)$ (6)
RND
Random network distillation (RND) was applied to provide an exploration bonus to each agent to universally novel molecules. This bonus is calculated as the squared error between a target network $f(x)$ and predictor network $\hat{f}(x)$ mapping of $x$. The predictor network is trained iteratively on all visited states $s\in\mathcal{S}$.
$R(\tau)^{\prime}=R(\tau)+\lambda\cdot||\hat{f}(x;\theta)-f(x)||^{2}$ (7)
The networks used in our implementation employ the same architecture as the CLM agents. The predictor network is trained to minimize the log-likelihood of SMILES sequences. The bonus is computed as the difference in policies between SMILES sequences measured by the summed log-likelihoods $\sum_{t=1}^{T}\log P_{\theta}(x_{t}|x_{<t})$.
ENTS
An additional loss term for the $k$-th agent was added to minimize the entropy across on-policy final states of agents $\mathcal{S}^{\{1,\dots,K\}\subseteq\mathcal{N}}$, to encourage the preferential sampling of states sampled by the $k$-th agent and discourage states sampled by other agents. To measure the likelihood of an agent sampling a state we used perplexity ($\mathcal{P}$).
$\begin{split}\mathcal{P}(\tau)&=e^{-\frac{1}{T}\sum_{t=1}^{T}\log\pi(a_{t}\mid s_{t})}\\ \end{split}$ (8)
The perplexities given the $k$-th agent policy $\pi_{k}$ were summed over on-policy states of agents $\mathcal{S}^{\{1,\dots,K\}\subseteq\mathcal{N}}$ resulting in a state-conditioned policy uncertainty value for each set of agent $i$ states $\mathcal{P}_{\text{sum}}(\pi_{k},\mathcal{S}^{i})$. The negative of this vector (so that smaller numbers represent more uncertainty) was softmax normalized resulting in a pseudo-probability $\tilde{P}(\mathcal{S}^{i}|\pi_{k})$ of agent $k$ sampling states derived from different agents.
$\begin{split}\mathcal{L}^{k}_{ENT_{\mathcal{S}}}&=\lambda\cdot\sum_{i=1}^{k}\tilde{P}(\mathcal{S}^{i}|\pi_{k})\log\tilde{P}(\mathcal{S}^{i}|\pi_{k})\end{split}$ (9)
CES
A natural extension from ENTS is the addition of a cross-entropy loss term. The same steps from ENTS were followed to result in a pseudo-probability $\tilde{P}(\mathcal{S}^{i}|\pi_{k})$ of agent $k$ sampling states derived from different agents. Then cross-entropy was applied to explicitly indicate that the $k$-th agent should assign a higher pseudo-probability to sampling states it has collected itself $\mathcal{S}^{k}$ than states collected by other agents $\mathcal{S}^{i}$.
$\begin{split}\mathcal{L}^{k}_{CE_{\mathcal{S}}}&=-\lambda\cdot\log\tilde{P}(\mathcal{S}^{k}|\pi_{k})\end{split}$ (10)
DIFFS
An additional loss term for the $k$-th agent to explicitly maximize the difference between uncertainty of states collected by previous agents relative to those collected by itself. Similar to ENTS the summed perplexity $\mathcal{P}_{\text{sum}}(\pi_{k},\mathcal{S}^{i})$ for collect states given the $k$-th agent policy was calculated.
$\mathcal{L}^{k}_{DIFF_{\mathcal{S}}}=-\lambda\cdot\lvert\frac{\sum_{i=1}^{k-1}\mathcal{P}_{\text{sum}}(\pi_{k},\mathcal{S}^{i})}{k-1}-\mathcal{P}_{\text{sum}}(\pi_{k},\mathcal{S}^{k})\rvert$ (11)
DIFFN
As initially proposed by Hu et al. (2024), an additional loss penalty was added which seeks to maximize the difference in log-probability for on-policy states $\mathcal{S}^{k}$ between the $k$-th agent and previous agents, scaled by the reward $R(\tau)$.
$\mathcal{L}^{k}_{DIFF_{\mathcal{N}}}=-\lambda\cdot\sum_{i=1}^{k-1}R(\tau)\cdot\lvert\log\pi_{k}(\tau)-\log\pi_{i}(\tau)\rvert$ (12)
DvD
Taking inspiration from Parker-Holder et al. (2020) we implemented Diversity via Determinant to enforce divergent behavior of agent policies. A random selection of 100 on-policy states were sampled and used to form an embedding that encoded policy behavior. This embedding was formed by concatenating the transition probabilities for possible actions $P(s^{i}_{t+1}|a^{i}_{t},s_{t})$ for a given agent policy $\pi^{i}_{\theta}$ for each sampled state $s$. Let this form the agent embedding $\mathbf{e_{i}}$ for each agent $i$ which were then used to compute pairwise similarities via a radial basis function (RBF) kernel:
$K_{ij}=\exp\left(-\frac{\|\mathbf{e}_{i}-\mathbf{e}_{j}\|^{2}}{2\sigma^{2}}\right),$ (13)
where $K_{ij}$ is the kernel value between agents $i$ and $j$, $\|\mathbf{e}_{i}-\mathbf{e}_{j}\|^{2}$ is the squared Euclidean distance, and $\sigma$ is a kernel bandwidth parameter.
The determinant of the kernel matrix $K$, denoted as $\det(K)$, was used as a measure of policy divergence which quantifies diversity, with higher values indicating greater divergence in agent policies and hence, was used as a penalty term:
$\mathcal{L}^{i}_{\text{DvD}}=-\det(K),$ (14)
POPNORM
Population normalization was used as a strategy to encourage divergent behavior from previous agents by adding a penalty term of the average population return up to the current $k$-th agent.
$\mathcal{L}^{k}_{\text{POPNORM}}=-\frac{1}{k-1}\sum^{k-1}_{i=1}\pi_{i}(\tau)\cdot R(\tau)$ (15)
Appendix C Molecular exploration benchmark
We selected four tasks, each with 2-4 objectives as molecular targets to be rediscovered based on real-world drug candidates ranging from pre-clinical (PC), clinical phases 1-3 (PH I-III) and approved (PH IV), shown in Figure 7 to Figure 10. For each task, the goal is to maximize the reward $R(m)$ for a given molecule $m$, where the reward is similarity to the closest molecular target $m_{t}$:
$R(m)=max(sim(m_{i},m_{t_{1}}),...,sim(m_{i},m_{t_{N}}))$ (16)
In this work, the similarity function $sim$ is implemented as the Levenshtein similarity between the canonical SMILES strings. This choice is because the CLM generates and acts in a SMILES environment. We created the Levenshtein similarity by normalizing the Levenshtein distance by the length of the reference string, clipped to ensure no negative values:
$sim_{\mathbb{L}}(m,m_{t})=clip(\frac{1-dis_{\mathbb{L}}(m,m_{t})}{|m_{t}|})$ (17)
Note that the SMILES strings of both $m$ $m_{t}$ are canonicalized and both used non-isomeric SMILES. Moreover,we made an alternative similarity function available using the Tanimoto similarity of ECFP fingerprints for application with non-language models, although the Levenshtein similarity will still work provided a SMILES representation is given.
Performance was evaluated based on the ability to generate molecules close to all molecular targets by taking the product:
$\prod_{t\in T}\frac{\sum_{i\in N}sim(m_{i},m_{t})}{N}$ (18)
This formulation tests the generative models ability to maximize the reward and remain intrinsically curious and continue exploring - even when the expected reward may already be high.
Figure 7: Antipsychotic (AP) drugs as molecular targets. Both are clinically approved antipsychotics, however, Lumateperone was approved in 1956, while Prochlorperazine was the most recently approved in 2019.
Figure 8: Adenosine A2A (A2A) receptor drug candidates as molecular targets. All the molecular candidates are A2A receptor antagonists that bind to the same orthosteric site but with alternative binding modes. XAC binds from N253 up through the pocket with a polar tail that extends into solvent, LUAA47070 binds via a water mediated interaction with N253, while Imaradenant binds from N253 down occupying lipophilic hotspots
Figure 9: Beta-secretase I (BACE1) receptor drug candidates as molecular targets. All are high-affinity inhibitors of BACE1 receptor with a common functional amidine/hiourea group to bind to the dual Asparctic acids in the pocket but are otherwise topologically different.
Figure 10: Epidermal growth factor (EGFR) drug candidates as molecular targets. All molecules are high-affinity covalent inhibitors of EGFR, sharing the reactive acrylamide functional group to bind to cysteine. Otherwise they are the four most topologically diverse molecules out of the 13 sub-100nM activity, at least phase I EGFR inhibitors found in ChEMBL34.
Appendix D Baseline algorithms
SCREEN: This virtual screening baseline iteratively samples without replacement from the CLM training dataset Appendix A curated from ChEMBL34. Note that all the target molecules from MolExp benchmark are included in the dataset, therefore, it is possible to achieve a perfect score if they are sampled. However, this is unlikely given the budget of 10,000 molecules out of the possible 1.7M.
REINFORCE: This baseline was implemented as the vanilla reinforce algorithm where the objective is to maximize the likelihood of sequences weighted by the cumulative reward $R(\tau)$. In this implementation we additionally applied experience replay to augment on-policy data with previously high rewarded molecules. All configuration parameters are shown in Table 2.
$\nabla J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\cdot R(\tau)\right]$ (19)
REINVENT: This baseline was implemented following the REINVENT formulation (Olivecrona et al., 2017) whereby a fixed version of the prior policy $\pi_{prior}$ (i.e., the pre-trained CLM) is used to regularize the updates of the agent policy $\pi_{agent}$ being trained. This is a form of reward shaping as shown in following Equation 20. This also introduces a hyperparameter $\sigma$ that determines the balance between $\pi_{prior}-\pi_{agent}$ difference or the reward $R(\tau)$. The hyperparameters used corresponding to the default frequently used in the literature are shown in Table 3.
$R(\tau)_{reshaped}=\frac{(\pi_{prior}-\pi_{agent}+\sigma\cdot R(\tau))^{2}}{\pi_{agent}}$ (20)
REINVENTMolOpt: This baseline is a hyperparameter optimized configuration of REINVENT for performance on the MolOpt benchmark. Notably this increases $\sigma$ to focus more on molecule reward higher than regularization to the prior and increases the amount of experience replay used. All configuration parameter are shown in Table 4.
[…]
Table 2: Hyperparameters for REINFORCE.
Hyperparameter Value
num_envs 128
total_smiles 10,000
model GRU (embedding of size 256 + 3 layer GRU of size 512 + MLP)
lr 0.0001
experience_replay True
replay_buffer_size 100
replay_batch_size 10
Table 3: Hyperparameters for REINVENT.
Hyperparameter Value
num_envs 128
total_smiles 10,000
model GRU (embedding of size 256 + 3 layer GRU of size 512 + MLP)
lr 0.0001
experience_replay True
replay_buffer_size 100
replay_batch_size 10
sigma 120
Table 4: Hyperparameters for REINVENTMolOpt.
Hyperparameter Value
num_envs 64
total_smiles 10,000
model GRU (embedding of size 256 + 3 layer GRU of size 512 + MLP)
lr 0.0005
experience_replay True
replay_buffer_size 100
replay_batch_size 24
sigma 500
Table 5: Hyperparameters for AHC.
Hyperparameter Value
num_envs 128
total_smiles 10,000
model GRU (embedding of size 256 + 3 layer GRU of size 512 + MLP)
lr 0.0001
experience_replay True
replay_buffer_size 100
replay_batch_size 10
sigma 60
topk 0.5
Table 6: Hyperparameters for ACEGENPractical.
Hyperparameter Value
num_envs 128
total_smiles 10,000
model GRU (embedding of size 256 + 3 layer GRU of size 512 + MLP)
lr 0.0001
experience_replay True
replay_batch_size 10
replay_buffer_size 100
replay_sampler uniform
sigma 0.005
topk 0.5
alpha 5
baseline mab
Table 7: Hyperparameters for ACEGENMolOpt.
Hyperparameter Value
num_envs 32
total_smiles 10,000
model GRU (embedding of size 256 + 3 layer GRU of size 512 + MLP)
lr 0.0001
experience_replay True
replay_batch_size 50
replay_buffer_size 100
replay_sampler prioritized
sigma 0.001
topk 0.5
alpha 3
baseline False
Appendix E Computational resource
All experiments detailed in this work were conducted on a single consumer grade GPU, more specifically an NVIDIA RTX 3090 with 1 CPU core.
Figure 11: Wall time required when running the MolExp benchmark with increasing total budget sizes (larger population size for multi-agent RL).
Appendix F Supplementary figures
Figure 12: Application of noise with varying coefficients $\lambda_{Noise}$ to either the final linear layer weights or all GRU hidden layers and the final linear layer weights. Increasing amounts of noise decrease the validity of molecules generated.
Figure 13: Baseline performance for each MolExp task during RL training, single replicate. Each line represents similarity to a single objective or target molecule. Note that ACEGEN and REINVENTMolOpt methods outperform due to their enhanced ability to optimize similarity to at-least one objective. Interestingly REINVENTMolOpt optimizes similarity to a different target molecule than ACEGEN for the EGFR task, despite identical initial policy parameterization.
Figure 14: Estimated maximum performance achieved by ACEGENMolOpt. ACEGENMolOpt was trained to maximize the similarity to each individual molecular target separately for a budget of 10,000 molecules.
Figure 15: Multiple independent agent learning for each MolExp task during RL training, single replicate. Each line represents average similarity to a single objective or target molecule, and for clarity the standard deviation is not plotted due to already high variance across agents. As the number of agents increases (increasing the total budget), by natural stochasticity it is more likely that there is divergence in agent behavior. Interestingly, even with 128 independent agents all seem to choose optimization of one molecule of the other for the AP task.
Figure 16: Single-agent learning for each MolExp task during RL training, single replicate. Each line represents similarity to a single objective or target molecule. In general, a single agent optimizes for similarity to one target molecule and doesn’t change throughout training. One notably exception exists for EGFR with a total budget of 640,000, where the agent seems to switch molecular target (in this case from green to blue).
Figure 17: Single-agent with RND exploration bonus for each MolExp task during RL training, single replicate. Each line represents similarity to a single objective or target molecule. In general, a single agent optimizes for similarity to one target molecule and doesn’t change throughout training, even with exploration bonuses. Note that drops in average score occur from a budget of 320,000 onward, but never this never changes the target molecule maximized.
Figure 18: Cooperative-agent learning for each MolExp task during RL training, single replicate, part I. Each line represents similarity to a single objective or target molecule. Training of each agent is plotted. Perfect cooperation is each agent learning to rediscover a different target molecule. Most cooperative strategies investigated result in no additional divergent behavior or in some cases, slower learning.
Figure 19: Cooperative-agent learning for each MolExp task during RL training, single replicate, part II. Each line represents similarity to a single objective or target molecule. Training of each agent is plotted. Perfect cooperation is each agent learning to rediscover a different target molecule. Most cooperative strategies investigated result in no additional divergent behavior or in some cases, slower learning.
Figure 20: Cooperative-agent learning for each MolExp task during RL training, single replicate, part III. Each line represents similarity to a single objective or target molecule. Training of each agent is plotted. Perfect cooperation is each agent learning to rediscover a different target molecule. Most cooperative strategies investigated result in no additional divergent behavior or in some cases, slower learning.

### v2

@@
-Test-Time Training Scaling for Chemical Exploration in Drug Design
+Test-Time Training Scaling Laws for Chemical Exploration in Drug Design
 Abstract
-Chemical language models for molecular design have the potential to find solutions to multi-parameter optimization problems in drug discovery via reinforcement learning (RL).
-A key requirement to achieve this is the capacity to ”search” chemical space to identify all molecules of interest.
-Here, we propose a challenging new benchmark to discover dissimilar molecules that possess similar bioactivity, a common scenario in drug discovery, but a hard problem to optimize.
-We show that a population of RL agents can solve the benchmark, while a single agent cannot.
-We also find that cooperative strategies are not significantly better than independent agents.
-Moreover, the performance on the benchmark scales log-linearly with the number of independent agents, showing a test-time training scaling law for chemical language models.
-Machine Learning, ICML
-[…]
+Chemical Language Models (CLMs) leveraging reinforcement learning (RL) have shown promise in de novo molecular design, yet often suffer from mode collapse, limiting their exploration capabilities.
+Inspired by Test-Time Training (TTT) in large language models, we propose scaling TTT for CLMs to enhance chemical space exploration.
+We introduce MolExp, a novel benchmark emphasizing the discovery of structurally diverse molecules with similar bioactivity, simulating real-world drug design challenges.
+Our results demonstrate that scaling TTT by increasing the number of independent RL agents follows a log-linear scaling law, significantly improving exploration efficiency as measured by MolExp.
+In contrast, increasing TTT training time yields diminishing returns, even with exploration bonuses.
+We further evaluate cooperative RL strategies to enhance exploration efficiency.
+These findings provide a scalable framework for generative molecular design, offering insights into optimizing AI-driven drug discovery.
 1 Introduction
-Identifying better, safer, and more efficacious drug candidates faster can save hundreds of millions of dollars (Bender & Cortés-Ciriano, 2021), and bring medicines to patients years quicker.
-However, the vastness of chemical space (Polishchuk et al., 2013) is a central challenge.
-It is intractable to computationally enumerate all drug-like molecules, let alone evaluate them, synthesize them and experimentally test them in the lab.
-Traditionally, virtual screening is conducted of virtual libraries that contain just tiny fractions of drug-like chemistry that can be purchased readily and experimentally tested.
-On the other hand, goal-directed generative AI can generate novel chemical structures with optimal predicted properties while implicitly searching much larger chemical spaces (Arús-Pous et al., 2019a).
-Thus, harnessing the full potential of generative AI to facilitate drug design will have considerable positive impacts on drug discovery as a whole.
-To fully exploit generative AI potential, achieving a thorough exploration of chemical space is critical.
-The more diverse solutions identified the more advantageous it is.
-Starting a drug discovery campaign from multiple diverse hits can offer backup series with alternative chemical and biological profiles, decreasing risk by increasing shots at goal (Brown & Boström, 2018).
-Additionally, probing novel areas of chemical space can avoid already patented compounds (Shimizu et al., 2023).
-Furthermore, in practice, many generative models undergo ”mode collapse” (a narrowing of chemical space) when optimizing for particular predicted properties, particularly those that are predicted by machine learning models with limited applicability domains arising from epistemic uncertainty (Preuer et al., 2018; Renz et al., 2019).
-Reinforcement learning (RL) is a family of optimization algorithms that can train an agent (a decision-making entity) to maximize a reward within a Markov decision process (Sutton & Barto, 2018).
-In generative molecular design, the generative model acts as the RL agent that learns to adapt molecule generation to maximize a given reward.
-One or more predicted molecular properties can be combined into an ’oracle’ that provides the reward.
-A range of generative models as RL agents for molecular design have been demonstrated including heuristic rules combined with deep Q-networks (Zhou et al., 2019), deep language model policy functions (Olivecrona et al., 2017; Ståhl et al., 2019), or deep graph network policy functions (Atance et al., 2022).
-However, it is the deep language model policy functions, also referred to as chemical language models (CLMs) (Segler et al., 2018; Grisoni, 2023; Skinnider et al., 2021), that have emerged as the most popular generative drug design algorithms (Olivecrona et al., 2017; Thomas et al., 2022b; Guo & Schwaller, 2024), as they have demonstrated top performance on multiple benchmarks (Brown et al., 2019; Polykovskiy et al., 2020; Thomas et al., 2022a).
-CLMs are autoregressive language models such as recurrent neural networks (RNNs) or GPT-style models trained on next token prediction typically using 1D SMILES (Weininger, 1988) strings that represent the 2D graph of a molecule.
-Seminal works (Olivecrona et al., 2017; Popova et al., 2018) applied RL to CLMs resulting in an agent capable of generating de novo molecules with desirable properties.
-The body of work now surrounding these models includes algorithmic variations to improve the learning efficiency (Bjerrum et al., 2023; Thomas et al., 2022b; Guo & Schwaller, 2024; Thomas et al., 2025), the practical utility (Fialková et al., 2021; Thomas et al., 2024a), as well as experimental validation of workflows utilizing such models for de novo drug design (Korshunova et al., 2022; Quinn et al., 2024; Thomas et al., 2024b).
-Interestingly, REINFORCE (Sutton & Barto, 2018) based RL algorithms remain superior for CLMs (Bou et al., 2024), likely due to strong policy initialization from pre-training (Ahmadian et al., 2024).
-In the field of Large Language Models (LLMS), Test-Time Training (TTT) has emerged as a successful approach to improve abstract reasoning (Akyürek et al., 2024).
-TTT is a strategy that updates the parameters of a pre-trained LLM on a specific test-time task, where they are reset to the pre-trained parameters for the next task.
-Hence, the focus is on the predictions of the temporarily updated LLM instead of the updated LLM with its new parameters.
-Moreover, scaling inference-time has shown to improve performance on benchmarks log-linearly with the number of samples generated at inference (Brown et al., 2024).
-Together, they have led to landmark improvements in abstract reasoning benchmarks such as the ARC Challenge (Chollet, 2019).
-Multi-agent reinforcement learning (MARL) involves the cooperation, competition, or evolution of multiple agents acting in an environment to enhance learning (Buşoniu et al., 2010).
-This is a proven effective strategy to reach human or greater level performance in StarCraft (Vinyals et al., 2019) and capture the flag (Jaderberg et al., 2019), or to learn cooperative microeconomic strategies (Johanson et al., 2022).
-Most relevant to the work presented here is the capacity of a population of agents to increase the diversity of solutions via cooperative learning by enforcing divergent behavior between agents (Parker-Holder et al., 2020).
-Although MARL is effective in fields such as game theory, few examples exist applying MARL to generative molecular design.
-In this work, we provide two central contributions:
-  1. 1.
-    We introduce a new hard de novo drug design benchmark focused on efficient Molecule Exploration (MolExp).
-Unlike existing benchmarks, MolExp emphasizes discovering structurally diverse molecules that retain similar bioactivity, acting as an exploration challenge where multiple distinct regions of chemical space get high rewards and all should be found.
-  2. 2.
-    We show that a population of independent agents is better at chemical space exploration on the benchmark than a single agent trained for a long time, even with exploration bonuses.
-We also show that performance on the benchmark scales with the number of agents in the population.
-Lastly, we investigate several cooperative strategies between multiple agents to try to enhance exploration efficiency.
+The vast chemical space, estimated to contain over $10^{60}$ drug-like molecules [1], presents a formidable challenge in drug discovery.
+Efficiently identifying diverse, safe, and efficacious drug candidates can accelerate development timelines and reduce costs by hundreds of millions of dollars [2].
+Traditional virtual screening of limited chemical libraries captures only a fraction of possible compounds, restricting exploration.
+In contrast, goal-directed generative AI, particularly Chemical Language Models (CLMs), enables de novo design of novel molecules with optimized properties, implicitly navigating expansive chemical spaces [3].
+Thorough exploration of chemical space is critical to maximize the potential of generative AI in drug design.
+Identifying structurally diverse molecules with similar bioactivity increases the likelihood of discovering backup series with distinct chemical and biological profiles, mitigating risks in drug development [4].
+Moreover, exploring novel chemical regions can circumvent patented compounds [5].
+However, many generative models, including CLMs, suffer from mode collapse when optimizing for predicted properties, particularly those reliant on machine learning models with limited applicability domains due to epistemic uncertainty [6, 7].
+CLMs, autoregressive models trained on SMILES strings [8], have become the leading approach for generative drug design [9, 10, 11], excelling in benchmarks like GuacaMol [12, 13].
+By integrating reinforcement learning (RL), seminal works [14, 15] have enabled CLMs to generate molecules with tailored properties.
+Advances in RL algorithms [16, 17, 18, 19] and practical applications [20, 21] have further enhanced their utility, with REINFORCE-based methods remaining superior due to robust pre-training [22, 23, 24].
+Inspired by Test-Time Training (TTT) in large language models, where model parameters are temporarily updated for specific tasks [25], we frame RL-based optimization of pre-trained CLMs as a form of TTT.
+In LLMs, scaling inference-time sampling has shown log-linear performance improvements [26].
+Similarly, scaling TTT for CLMs—by increasing the number of RL agents or training time—offers potential to enhance chemical space exploration.
+Cooperative RL strategies further amplify exploration through population-based learning [27], though their application in molecular design remains underexplored.
+In this work, we make two key contributions.
+First, we demonstrate that scaling TTT by increasing the number of independent RL agents follows a log-linear scaling law, significantly enhancing exploration efficiency.
+Conversely, extending TTT training time yields limited benefits, even with exploration bonuses or diversity filters.
+We also propose and evaluate cooperative RL strategies to further improve exploration efficiency.
+Second, we introduce MolExp, a novel benchmark to measure Molecule Exploration.
+MolExp requires the rediscovery of structurally diverse molecules with comparable bioactivity.
+Unlike existing benchmarks, MolExp requires exploration of all high-reward regions of chemical space, extending the boundaries of real-world drug design.
+This work provides a robust framework for enhancing chemical space exploration, addressing critical challenges in AI-driven drug discovery.
+By leveraging TTT and cooperative RL, we offer practical tools through MolExp and scalable strategies that can accelerate the discovery of diverse drug candidates.
 2 Related Work
-Balancing the exploration-exploitation trade-off is a constant challenge in RL, including RL for drug design (Langevin et al., 2024).
-Exploitation initially maximizes the reward quicker but may lead to a lower overall reward, while exploration may lead to higher overall reward and increased diversity of solutions but takes longer.
-Diversity has consistently been an important topic in generative molecular design and is increasingly included in the evaluation of performance (Thomas et al., 2022a; Renz et al., 2024).
-Several approaches have been proposed to increase diversity, for example, using dual networks to balance exploitation-exploration during sampling (Liu et al., 2019), penalizing the reward of molecules similar to previously generated ones (Blaschke et al., 2020), or increasing the reward of molecules that are dissimilar to those previously generated (Thiede et al., 2022).
-Recent work shows that random network distillation (RND) (Burda et al., 2018) is particularly effective at increasing molecular diversity with low overhead due to not requiring an explicit memory of previously generated molecules (Svensson et al., 2024; Thomas et al., 2025).
-Previous work has applied multi-agent RL to de novo drug design.
+Balancing the exploration-exploitation trade-off is a constant challenge in RL, including RL for drug design [28].
+Exploitation initially maximizes the reward quicker but may lead to a lower overall reward, while exploration may lead to a higher overall reward and increased diversity of solutions but takes longer.
+Diversity has consistently been an important topic in generative molecular design and is increasingly included in the evaluation of performance [29, 30].
+Several approaches have been proposed to increase diversity, for example, using dual networks to balance exploitation-exploration during sampling [31], penalizing the reward of molecules similar to previously generated ones [32], or increasing the reward of molecules that are dissimilar to those previously generated [33].
+Recent work shows that random network distillation (RND) [34] is particularly effective at increasing molecular diversity with low overhead due to not requiring an explicit memory of previously generated molecules [35, 19].
+Previous work has applied population-based RL to de novo drug design.
 Hu et al.
-(2024) deployed up to 4 GPT agents as CLMs that use the REINVENT loss formulation with an additional penalty to discourage the $k$th agent from being similar to previous agents.
-This multi-agent setup outperformed baseline algorithms on the GuacaMol benchmark (Brown et al., 2019).
+[36] deployed up to 4 GPT agents as CLMs that use the REINVENT loss formulation with an additional penalty to discourage the $k$-th agent from being similar to previous agents.
+This multi-agent setup outperformed baseline algorithms on the GuacaMol benchmark [12].
 However, this benchmark consists of mostly simple tasks (as also stated by the authors).
-As demonstrated by their experiment comparing 1, 2, 4 and 8 agents all having near maximum performance and all within error.
+As demonstrated by their experiment comparing 1, 2, 4 and 8 agents all having near maximum performance and all within standard deviation.
 Moreover, independent agents and other cooperative strategies were not investigated.
-Therefore it remains to be seen whether MARL improves performance significantly, by how much, and which cooperative strategies work best.
+Therefore it remains to be seen whether cooperative strategies improve performance significantly, by how much, and which cooperative strategies work best.
 Generative molecular design benchmarks must reflect realistic scenarios in drug design for practical application, as gold standard evaluation via synthesis and experimental validation of molecules is too costly at scale.
-The GuacaMol benchmark (Brown et al., 2019) proposed a suite of 20 tasks, the authors highlighted that $\sim 15/20$ tasks are easily solved by the generative algorithms.
+The GuacaMol benchmark [12] proposed a suite of 20 tasks, the authors highlighted that $\sim 15/20$ tasks are easily solved by the generative algorithms.
 The MolOpt benchmark limited these tasks to a budget of 10,000 molecules during optimization, shifting the focus to efficiency, however, this likely over-rewards exploitative algorithms as opposed to explorative ones.
-New criteria have been defined to additionally measure the quality of chemistry generated which has changed rank performance (Thomas et al., 2022a; Renz et al., 2024), moreover, rank performance on this MolOpt benchmark didn’t reflect performance on more realistic optimization objectives such as predicted binding affinity via docking (Bou et al., 2024).
+New criteria have been defined to additionally measure the quality of chemistry generated which has changed rank performance [29, 30], moreover, rank performance on this MolOpt benchmark didn’t reflect performance on more realistic optimization objectives such as predicted binding affinity via docking [23].
 Hence, there remains significant scope for new exploration-focused objectives that better reflect realistic optimization objectives in drug design.
@@
 The scoring function (also referred to as an oracle) $s(x)\in[0,1]$ may combine one or more property predictors $p_{i}(x)$ that score a given molecule $x$ returning a real value $p_{i}:X\rightarrow\mathbb{R}$.
-Multi-parameter optimization can be achieved by combining multiple predicted properties, the simplest example being to take the arithmetic mean $\sum_{i}t_{i}(p_{i}(x))$ ensuring that $t(p_{i}(x))\in[0,1]$ by using a transformation function $t_{i}$.
-Overall, the final score reflects the desirability of the molecule for the given context, however, optimizing a score combining many properties is usually non-trivial as improving one property can be detrimental to another resulting in a Pareto-optimization problem.
-$\max_{x\in X}s(x)$ (1)
-[…]
+The final score reflects the desirability of the molecule for the given context, where the goal is to generate a molecule with the best score.
+$\arg\max_{x\in X}s(x)$ (1)
+However, chemical space is vast, property predictors are inherently inaccurate with high uncertainty, and Pareto optimal solutions have different property profiles.
+Hence, multiple molecules exist across chemical space that are considered hits if the score is above a given threshold $T$ (or sometimes more simply the top-$k$ molecules).
+We propose that the problem of generative molecular design should extend to finding all possible hits across chemical space, and postulate that this better aligns with practical requirements in drug discovery.
 $\exists x_{i},...,x_{N}\in X,\,x_{i}\neq x_{j}\text{ such that }s(x_{i})\geq T$ (2)
-We propose that a generative model benchmark should adequately mimic this concept of identifying multiple, seemingly unrelated molecular targets in chemical space (Figure 1, right).
+We propose that a generative model benchmark should adequately mimic this concept of identifying multiple, even structurally unrelated molecular targets in chemical space (Figure 1, right).
 Firstly, previously published benchmarks typically don’t capture this with objectives focused on a single narrow target, or a single broad target (Figure 1, left).
 These do not adequately test the ability of an algorithm to explore chemical space if it is already receiving high rewards.
-Secondly, there are post-hoc metrics that measure the diversity of solutions (Thomas et al., 2022a; Renz et al., 2024), however, this only measures the diversity inherently embedded in the objective.
+Secondly, there are post-hoc metrics that measure the diversity of solutions [29, 30], however, this only measures the diversity inherently embedded in the objective.
 Therefore, designing an objective to contain multiple distinct targets provides a greater opportunity to observe inherent exploration capability.
-Lastly, previous benchmarks predominantly measure performance on the maximum $S(x)$ achieved, whereas, we propose to measure performance on maximum $p_{i}(x)$ for each molecular target.
+Lastly, previous benchmarks predominantly measure performance on the maximum $S(x)$ achieved, whereas, we propose to measure performance on maximum $p_{i}(x)$ for each molecular target representing a maxima in the reward landscape.
 This focuses evaluation on the algorithm’s ability to explore all high-rewarding regions.
@@
 3.2 Reinforcement learning with chemical language models
-CLMs utilize a language representation, commonly SMILES (Weininger, 1988), to leverage natural language models such as recurrent neural networks (RNNs) or GPT-style models.
-For example, SMILES represent the molecular structure of paracetamol with the string ”CC(=O)Nc1ccc(O)cc1”.
+CLMs utilize a language representation, commonly SMILES [8], to leverage natural language models such as recurrent neural networks (RNNs) or GPT-style models.
+For example, SMILES represent the molecular structure of paracetamol with the string “CC(=O)Nc1ccc(O)cc1".
 These models operate autoregressively, predicting the next token $x_{t}$ in a sequence $X=(x_{1},x_{2},\ldots,x_{T})$ based on the context provided by previous tokens.
@@
 We deviate to vanilla REINFORCE by reshaping the reward to regularize the agent policy to the prior policy $\pi_{prior}$ with a coefficient $\sigma$ and by adding an exponent $\alpha$, as proposed by Thomas et al.
-(2025).
+[19].
 These components improve the quality of molecules proposed by the agent and improve performance by steepening the gradients in the reshaped reward landscape.
@@
 A replay buffer is kept of high-reward molecules which is sampled with prioritization proportional to their reward.
-This additionally improves learning efficiency (Gao et al., 2022; Thomas et al., 2025).
+This additionally improves learning efficiency [37, 19].
 $R(\tau)_{reshaped}=clip(R(\tau)+\sigma\cdot\pi_{prior})^{\alpha}$ (4)
-3.3 Multi-agent reinforcement learning
-MARL extends the framework of single-agent RL to systems with multiple independent or interacting agents, each learning to optimize its own policy in an environment.
+3.3 Cooperative reinforcement learning
+Cooperative RL extends the framework of single-agent RL to systems with multiple independent or interacting agents, each learning to optimize its own policy in an environment.
 The problem is the same as defined before, but now with $\mathcal{N}$ agents.
@@
 Note that independent agents will still deviate due to the stochastic nature of the transition probabilities $P(s^{i}_{t+1}|a^{i}_{t},s^{i}_{t})$ leading to the sampling of different actions $a^{i}_{t}$.
-In the interaction case, we focused on methods of cooperation within a population-learning setting which typically results in more diverse solutions (Oroojlooy & Hajinezhad, 2023).
+In the interaction case, we focused on methods of cooperation within a population-learning setting which typically results in more diverse solutions [38].
 In this case, using $\langle\mathcal{S},\mathcal{A}^{i},R,P^{i}\rangle$ for $i\in\mathcal{N}$ where the states $S$ and $R$ may be shared across agents in a cooperative manner.
@@
 DIFFN
-Re-implementation of the cooperative strategy applied by Hu et al. (2024) which seeks to maximize the log-probability difference of on-policy states between the current agent and other agents.
+Re-implementation of the cooperative strategy applied by Hu et al. [36] which seeks to maximize the log-probability difference of on-policy states between the current agent and other agents.
 DvD
-Diversity via Determinant (DvD) inspired by Parker Holder et al.
-(2020).
+Diversity via Determinant (DvD) inspired by Parker-Holder et al.
+[27].
 Calculate a policy behavior embedding and maximize the divergence in policy behavior as measured by the determinant of the behavioral embedding transformed by a kernel function.
@@
 4 Experiments
-To assess the ability of single and multiple RL agents to explore chemical space, we conducted three experiments.
+First, we describe our proposed benchmark MolExp, and show performance of various baseline RL algorithms with a comparison to an established benchmark.
+Then, we conducted three experiments to assess chemical exploration.
   1. 1.
-    We defined a molecule exploration benchmark and evaluated baseline algorithm performance.
+    We investigated the effect of scaling TTT by the number of RL agents, or scaling TTT by the training time of a single RL agent by its molecule budget.
   2. 2.
-    We tested the effectiveness of different population sizes of multiple independent agents, and compared them to equivalent single agents.
+    We proposed and tested different cooperative RL strategies to enhance exploration efficiency.
   3. 3.
-    We investigated different strategies of cooperation within a population of agents to enhance exploration efficiency.
+    We translated our findings to a practically relevant task in optimizing the predicted bioactivity of generated compounds.
 4.1 Molecule exploration benchmark
-To mimic our proposed problem formulation of identifying all hits in chemical space requiring intrinsic exploration, we designed a new molecular exploration (MolExp) benchmark.
-To embed this benchmark in real-world discovery, it was designed based on the converse of the similarity principle  [^1] In drug discovery, the similarity principle states that similar molecules $Sim(m_{1},m_{2})\geq\epsilon$ possess similar bioactivity $|f(m_{1})-f(m_{2})|\leq\delta$.
-However, this does not prove the converse and many exceptions exist that are more difficult to model., i.e., to identify multiple dissimilar molecules that exhibit similar desirable properties such as bioactivity.
-Note that activity cliffs are another example converse to the similarity principle that have been used to demonstrate more challenging scenarios for predictive models (Kwapien et al., 2022; Van Tilborg et al., 2022).
-Each run is replicated on the benchmark 5 times and as with previous benchmarks (Gao et al., 2022) a default budget of 10,000 molecules is used.
-Oracles The benchmark consists of four tasks, each using an oracle to evaluate molecules.
-These are based on drug candidates and their biological targets, including anti-psychotic drugs (AP), Adenosine A2A (A2A) receptor ligands, beta-secretase 1 (BACE1) inhibitors, and epidermal growth factor inhibitors (EGFR).
-These tasks contain 2, 3, 3, & 4 objectives respectively, where the objective is the similarity to a target molecule $sim(m_{i},m_{t})$.
-For each task, only a single reward is provided based on the similarity to the nearest target molecule $R(m)=max(sim(m_{i},m_{t_{1}}),...,sim(m_{i},m_{t_{N}}))$.
-Due to CLMs acting in string space, we chose a string-edit similarity function.
-This, reflects chemical space search scenarios where distinct regions of reward signal exist.
+To address our proposed problem formulation to identify all hits in chemical space, we designed a new molecular exploration (MolExp) benchmark.
+Crucially, a ground truth is required to know that all solutions are identified  [^1] Oracles that are more practically relevant, such as bioactivity prediction or docking, don’t contain a ground truth set of all possible high scoring solutions..
+For this reason, we use the toy-task of multiple molecule rediscovery via optimization of a similarity scoring function.
+To ensure relevance to real-world discovery, each set of molecules are at minimum pre-clinical drug candidates bioactive against the same target.
+Moreover, to make the objective more challenging, the set of molecules are selected to be structurally dissimilar albeit possessing similar bioactivity.
+This phenomena is the converse of the similarity principle  [^2] In drug discovery, the similarity principle states that similar molecules $Sim(m_{1},m_{2})\geq\epsilon$ possess similar bioactivity $|f(m_{1})-f(m_{2})|\leq\delta$.
+However, this does not prove the converse and many exceptions exist that are more difficult to model..
+Note that activity cliffs are another example converse to the similarity principle that have been used to demonstrate more challenging scenarios for predictive models [39, 40].
+Each run is replicated on the benchmark 5 times with a default budget of 10,000 molecules per run, consistent with previous benchmarks [37].
+Oracles
+The benchmark consists of four tasks.
+Each task is based on a set of drug candidates and their shared biological target, including anti-psychotic drugs (AP), Adenosine A2A (A2A) receptor drug candidates, beta-secretase 1 (BACE1) drug candidates, and epidermal growth factor drug candidates (EGFR).
+These tasks contain a set of 2, 3, 3, & 4 target molecules respectively, where the objective is to rediscover the full set of target molecules.
+To mimic a chemical space where distinct regions of reward signal exist, the oracle returns the similarity to the nearest target molecule as the reward $R(m)=max(sim(m_{i},m_{t_{1}}),...,sim(m_{i},m_{t_{N}}))$.
+As CLMs act in string space, we chose a string-edit similarity function based on Levenshtein distance (MolExpL).
+However, we additionally report the results when using fingerprint similarity (MolExp) in the Appendix.
 A more detailed description of the molecular targets chosen and the similarity function can be found in Appendix C.
-Metric Performance is measured by calculating the product of the top N most similar molecules to each target molecule respectively $\prod_{t\in T}\frac{\sum_{i\in N}sim(m_{i},m_{t})}{N}$.
+Metric
+Performance is measured by calculating the product of the top N most similar molecules to each target molecule respectively $\prod_{t\in T}\frac{\sum_{i\in N}sim(m_{i},m_{t})}{N}$.
 This metric focuses on the ability generate similar molecules to all objectives.
-Unless stated otherwise, we use $N=1$.
-Baselines We tested virtual screening and multiple RL algorithms (see Appendix D) on their inherent ability for chemical exploration on the MolExp benchmark.
-Table 1 shows that the recently published ACEGENMolOpt (Thomas et al., 2025), performed best.
-Figure 13 illustrates that none of the tested algorithms optimize all molecular targets within the default budget (for a single replicate).
-Moreover, that the best scores are due to improved efficiency in maximizing similarity to a single target molecule.
+Unless stated otherwise, we use $N=1$ reflecting the maximum similarity achieved.
+Molecular diversity is also measured by sphere exclusion diversity (SEDiv) at a sample size of 1,000 [41] which provides a more accurate representation of chemical space coverage than internal diversity [42, 30].
+Baseline performance
+We tested virtual screening, MolRL-MGPT [36], and multiple ACEGEN RL algorithms [23, 19] on their inherent ability for chemical exploration on the MolExpL benchmark (see Appendix D for implementation details).
+Note that MolRL-GPT was pre-trained on the same dataset for standardized comparison (see Appendix A for details).
+Table 1 shows that with a budget of 10,000 the ACEGENMolOpt algorithm achieved the highest score, therefore, we use this RL configuration for all further experiments.
+However, Figure 13 illustrates that none of the tested algorithms optimize all molecular targets within the default budget (for a single replicate), and that the best scores are due to improved efficiency in maximizing similarity to a single target molecule.
 Therefore, there is still considerable room for improvement.
-Estimated max An estimated maximum performance can be calculated for the chosen agent RL configuration (ACEGENMolOpt) by training the agent to maximize the similarity to each individual molecular target with the reward $R(m)=sim(m_{i},m_{t})$ only, Figure 14.
-This also tests the agents ability to optimize for the similarity to each individual molecular target in the first place which is an expected prerequisite for performance on the benchmark.
-Averaging the Top-1 similarity achieved over the 5 replicates and calculating the product over the group of molecular target objectives in each task, and then summing over all tasks led to an estimated achievable maximum benchmark score of 2.91.
-Table 1: Baseline performance of virtual screening and policy-gradient based RL algorithms on the MolExp benchmark within a budget of 10,000 molecules.
-ACEGENMolOpt outperforms all other methods, while virtual screening outperforms two baselines algorithms vanilla REINFORCE and REINVENT.
-Task SCREEN REINFORCE REINVENT REINVENTMolOpt AHC ACEGENPractical ACEGENMolOpt
-AP 0.53 ± 0.04 0.56 ± 0.05 0.57 ± 0.04 0.63 ± 0.02 0.62 ± 0.06 0.64 ± 0.02 0.64 ± 0.04
-A2A 0.38 ± 0.03 0.36 ± 0.02 0.39 ± 0.08 0.38 ± 0.03 0.40 ± 0.02 0.40 ± 0.07 0.44 ± 0.06
-BACE1 0.29 ± 0.08 0.23 ± 0.02 0.21 ± 0.01 0.27 ± 0.03 0.23 ± 0.01 0.23 ± 0.03 0.28 ± 0.03
-EGFR 0.17 ± 0.04 0.15 ± 0.02 0.16 ± 0.02 0.22 ± 0.03 0.17 ± 0.03 0.17 ± 0.02 0.25 ± 0.05
-Sum 1.36 ± 0.10 1.30 ± 0.06 1.33 ± 0.09 1.50 ± 0.06 1.41 ± 0.07 1.44 ± 0.08 1.62 ± 0.09
-4.2 Test-time training with RL for chemical exploration
-We trained a population of independent agents using the best-performing baseline (ACEGENMolOpt) configuration.
-In this case, each agent had a budget of 10,000 molecules.
-The results of different population sizes up to 128 agents is shown in Figure 2.
-We found that the benchmark performance continued to increase log-linearly from $\sim 1.6$ to $\sim 3.5$ up until the maximum population size was tested, almost reaching the maximum of score of 4.
-Interestingly, when the population size reached 32, the estimated maximum score for this RL configuration was exceeded.
-This is indicative of the power of leveraging the variance between different agents to overcome the more ambiguous reward signal, which did not turn out to be a limiting factor.
-Figure 15 shows the average similarity to each target during training across multiple agents, indicating the variance observed between multiple agents with different agents maximizing similarity to different molecular targets, resulting in overall increased exploration and increased benchmark scores.
-This is exemplified in Figure 3 which shows that a population of 128 agents identifies all 3 A2A targets.
-Interestingly, for all population sizes in the AP task, all agents maximized the same molecular target, suggesting a particular difficulty in optimizing the other.
-Figure 2: Benchmark performance on MolExp as the number of independent agents increases.
-The score reported is the product of the maximum similarity to each target molecule summed over the tasks.
-Note that performance increase scales log-linearly with increasing independent agents.
-Estimated max is calculated based on a single agent optimizing each task’s molecular target individually.
-Figure 3: Example molecules generated during the A2A task from the MolExp benchmark.
-In columns, the molecule with the maximum similarity to each target is shown, alongside which agent sampled it.
-1 agent almost solves target 3, 4 agents solve target 3 and almost target 2, meanwhile 128 agents solve all 3 targets.
-Increasing the number of agents, each with an individual budget, effectively increases the total budget.
-Therefore, we tested the single agent with the cumulative budget of a population of agents, for comparison.
-Figure 4 shows that a single agent trained for the equivalent budget does not increase performance similarly.
-In this case, scaling the duration of training for a single agent results in a saturation of benchmark performance after approximately 40,000 molecules evaluated, as agents expectedly fail to continue exploring once a target has been maximized (Figure 16).
-This is only improved slightly if state-of-the-art exploration strategies are employed such as RND, the most performant exploration strategy (Svensson et al., 2024; Thomas et al., 2025), which provides an exploration bonus to the reward.
-Where Figure 17 shows only small fluctuations in similarity due to RND, before subsequent maximization again to the same ’chosen’ molecular target.
-This is instead of maximization of similarity to a given molecular target and then a switch from one to another, as intended with exploration methods.
-Figure 4: Performance comparison between using multiple independent agents or running one single agent for longer.
-Note that running multiple independent agents with a budget of 10,000 each outperforms running a single agent with a combined budget of $10,000\times N_{agents}$.
-Additional exploration strategies such as RND don’t improve performance to the level of multiple independent agents.
-4.3 Cooperative multi-agent RL for chemical exploration
-Cooperation within a population of RL agents can theoretically improve performance by leveraging useful information shared between agents, such as, states already visited and their corresponding reward (Oroojlooy & Hajinezhad, 2023).
+These results are also observed when using fingerprint similarity oracles (see subsection E.2).
+Estimated max
+An estimated maximum performance can be calculated for ACEGENMolOpt RL configuration by training the agent to maximize only the similarity to each individual molecular target with the reward $R(m)=sim(m_{i},m_{t})$, (see Figure 14).
+This tests the agents ability to maximize similarity to each individual molecular target only, an expected prerequisite for maximizing similarity to multiple target molecules in the benchmark.
+Calculating the equivalent benchmark score results in an estimated maximum MolExpL score of 2.91.
+Comparison to GuacaMol
+For reference, baseline algorithms were tested on the GuacaMol benchmark also with a fixed budget of 10,000 molecules.
+This shows ACEGENMolOpt achieving the highest score (see subsection E.3).
+However, a key difference is that 5/6 GuacaMol similarity and rediscovery tasks are already solved by ACEGENMolOpt with a score $\geq 0.9$, whilst the MolExpL benchmark results in lower scores reflecting a higher difficulty of benchmark tasks.
+Table 1: Baseline performance on the MolExpL benchmark. The ACEGENMolOpt RL configuration performs best in 3/4 tasks which we carry forward for scaling experiments.
+Task SCREEN MolRL-MGPT REINFORCE REINVENT REINVENTMolOpt AHC ACEGENPractical ACEGENMolOpt
+AP 0.53 ± 0.04 0.48 ± 0.03 0.56 ± 0.05 0.57 ± 0.04 0.63 ± 0.02 0.62 ± 0.06 0.64 ± 0.02 0.64 ± 0.04
+A2A 0.38 ± 0.03 0.37 ± 0.03 0.36 ± 0.02 0.39 ± 0.08 0.38 ± 0.03 0.40 ± 0.02 0.40 ± 0.07 0.44 ± 0.06
+BACE1 0.29 ± 0.08 0.20 ± 0.02 0.23 ± 0.02 0.21 ± 0.01 0.27 ± 0.03 0.23 ± 0.01 0.23 ± 0.03 0.28 ± 0.03
+EGFR 0.17 ± 0.04 0.17 ± 0.01 0.15 ± 0.02 0.16 ± 0.02 0.22 ± 0.03 0.17 ± 0.03 0.17 ± 0.02 0.25 ± 0.05
+Sum 1.36 ± 0.10 1.21 ± 0.05 1.30 ± 0.06 1.33 ± 0.09 1.50 ± 0.06 1.41 ± 0.07 1.44 ± 0.08 1.62 ± 0.09
+4.2 Test-time training scaling for chemical exploration
+We tested if scaling TTT with larger population sizes of independent RL agents would increase chemical exploration and improve benchmark performance.
+In this case, we chose the best-performing RL configuration-ACEGENMolOpt-and assigned a budget of 10,000 molecules to each agent in the population.
+2(a) shows that MolExpL performance increases log-linearly from $\sim 1.6$ with 1 agent to $\sim 3.5$ with 128 agents, almost reaching a maximum score of 4.
+Log-linear scaling was also confirmed using the vanilla REINFORCE RL configuration, albeit with lower benchmark scores.
+Interestingly, when the population size reached 32 the estimated maximum score for this RL configuration was exceeded, indicating that the additional variance between independent agents overcomes the additionally reward signal ambiguity introduced from multiple molecular targets.
+Figure 17 shows that by increasing agents, there is more variance across the set of molecule targets.
+The AP task is an exception, where all 128 independent agents all maximize similarity to the same target molecule in the set.
+Similar results are also observed when using fingerprint similarity oracles subsection E.2.
+Additionally, similar scaling was achieved with up to 32 agents on the GuacaMol benchmark (subsection F.3), at which point performance on many tasks becomes saturated.
+(a) Scaling RL agents
+(b) Scaling budget
+(c) Molecular diversity
+Figure 2: Performance on the MolExpL benchmark with scaling.
+(a) Scaling the number of independent ACEGENMolOpt and REINFORCE RL agents, each with a budget of 10,000.
+(b) Scaling the total budget allocated to a single agent, a single agent with a RND exploration bonus, and a single agent with a DF.
+(c) The diversity of sampled compounds as measured by sphere exclusion diversity.
+Increasing the number of agents, each with a budget of 10,000, increases the total budget.
+Therefore, we compared the alternative approach of scaling TTT via increasing the budget for a single RL agent.
+However, this did not result in similar MolExpL performance gain as shown in 2(b).
+Single RL agent performance saturates at a budget of 40,000 as the agent expectedly fails to continue exploring once one of the target molecules has been maximized (Figure 18).
+Moreover, 2(c) shows that molecular diversity decreases while scaling budget.
+Employing RND as a state-of-the-art exploration strategy, the most performant exploration strategy [35, 19], marginally improves performance and molecular diversity.
+Where Figure 19 shows that RND only results in fluctuations in similarity to one molecular target, as opposed to the desired behavior of switching from one molecular target in the set to another.
+We additionally tested a simple diversity filter (DF) that assigns a reward of 0 to non-unique canonical SMILES.
+However, the DF also failed to rescue TTT scaling in this dimension.
+Figure 3 exemplifies how 87 agents are required to solve the A2A task and recover all target molecules, while scaling one agent results in a focus only on the third target molecule.
+Figure 3: Example molecules generated compared to the set of targets in the MolExpL A2A task.
+The maximum similarity score and corresponding agent $k$ is labeled.
+Note that only with 87 agents is the task solved.
+4.3 Cooperative multi-agent RL for targeted chemical exploration
+Cooperation within a population of RL agents can theoretically improve performance by leveraging useful information shared between agents, such as, states already visited and their corresponding reward [38].
 We explored several strategies to leverage cooperative learning to improve the efficiency of chemical exploration.
-Considering the maximum number of task objectives in the benchmark was 4, we investigated strategies utilizing 4 RL agents.
-Where, the best-case scenario is if each agent learns a different objective.
-The MolExp score of the different cooperative strategies are shown in Figure 5.
-All of the strategies tested were similar or worse in performance, with no statistically significant improvement over independent agents.
-The best strategies were POPNORM, ENTS, and DIFFN.
-Where lower $\lambda$ values performed best for ENTS showing divergent behavior for a single replicate in Figure 18 with little restriction in learning capability, although this is observed at larger values of $\lambda$.
-Meanwhile either the lowest (0.001) or highest (1) tested $\lambda$ value for DIFFN performed well.
-Interestingly, there didn’t seem to be any obvious correlation between the $\lambda$ value and restriction on learning capability for the values tested (Figure 20).
-Note that the best values here differ from the default of 0.01 used by Hu et al.
-(2024), also used with 4 CLM agents.
-The Noise strategy exhibited higher standard deviation in benchmark score.
-Both Purge and Shared replay buffer strategies dropped performance, with Shared being among the worst performing benchmarks as anticipated.
-Where in the example replicate shown in Figure 18 every agent optimizes the same molecular target and hence converged in chemical space as anticipated, with the novelty bonus failing to ameliorate this effect.
-Purge may not offer strict enough segregation of replay buffers to diverge agent behavior, as unique but similar molecules may still exist in different replay buffers.
-On the other hand, strategies like RND and DvD that should more strongly encourage divergent behavior appear to ultimately restrict learning within the defined budget (Figure 18), without resulting in the desired outcome of different agents optimizing different molecular targets more than independent agents.
-Lastly, no tested values of $\lambda$ for CES or DIFFS outperformed independent agents.
-Where higher values also led to restricted learning capability, especially in the case of DIFFS as shown in Figure 19 and Figure 20.
-Figure 5: Performance comparison of different multi-agent cooperative strategies on the MolExp benchmark.
-Each plot is a different cooperative strategy using 4 agents.
-The dashed line represents the average performance of 4 independent agents.
-Note that no strategy tested significantly outperformed a population of 4 independent agents as measured by a Bonferroni-corrected, one-tailed t-test.
+Considering the maximum number of molecular targets in a task set was 4, we investigated strategies utilizing 4 RL agents.
+Where, the best-case scenario is each agent maximizing similarity to a different molecular target in the set.
+We investigated an extensive range of the different cooperative strategies on the MolExpL benchmark, including the MolRL-MGPT baseline with the same budget and number of agents, shown in 4(a).
+All of the strategies tested were similar or worse in performance, with the best being POPNORM, ENTS, and DIFFN.
+However, no strategy significantly outperformed independent agents as measured by a Bonferroni-corrected, one-tailed t-test.
+On the other hand, 4(b) shows that many cooperative strategies successfully resulted in corresponding increases in molecular diversity, namely RND, ENTS, CES, and DIFFS.
+This is usually anti-correlated with MolExpL score.
+This highlights the difficulty and complexity in achieving targeted exploration, where diversity appears to hinder learning, or is naive rather than targeted on on favorable regions of chemical space.
+These trends were even more prominent when using fingerprint similarity oracles subsection G.2.
+Assessing the example behavior of different agents shows that some cooperative strategies do result in divergent behavior, but to the detriment of learning efficiency by slowing or inhibiting the learning of subsequent agents in a population (see subsection G.1).
+(a) Cooperative RL MolExpL benchmark performance
+(b) Cooperative RL molecular diversity
+Figure 4: Performance comparison of different 4-agent cooperative strategies on the MolExpL benchmark, each with a budget of 10,000.
+The dashed line represents the average of 4 independent agents as baseline.
+4.4 Maximizing predicted A2A bioactivity
+To represent a practically applicable objective function, we additionally tested TTT scaling behavior when maximizing the predicted probability of A2A bioactivity as measured by the QSAR classification model from MolScore [43].
+Increasing chemical exploration during optimization should increase the probability of recovering the set of known A2A drug candidates used in the MolExp A2A task, two of which are present in the training dataset positive class.
+Therefore, we use the same metric to measure the maximimum similarity achieved to the A2A target molecule set, we refer to this additional experiment as MolExpBio.
+However, this now depends on the properties of the QSAR model and therefore, is not theoretically guaranteed.
+5(a) shows that scaling the number of independent RL agents increases the benchmark score, again more than scaling the budget of a single RL agent (5(b)).
+In other words, scaling RL agents increases the maximum similarity to the set of target molecules, and hence the chance of rediscovering the drug candidates.
+This represents increased chemical exploration of the reward landscape.
+This task highlights the translational benefit of TTT scaling with RL agents to real-world drug design objectives.
+(a) Scaling RL agents
+(b) Scaling budget
+(c) Molecular diversity
+Figure 5: Performance on the MolExpBio task with scaling.
+(a) Scaling the number of independent ACEGENMolOpt and REINFORCE RL agents, each with a budget of 10,000.
+(b) Scaling the total budget allocated to a single agent, a single agent with a RND exploration bonus, and a single agent with a DF.
+(c) The diversity of sampled compounds as measured by sphere exclusion diversity.
 5 Conclusion
-In this paper, we present the challenge of exploring chemical space with a CLM to identify all possible candidate molecules in the context of drug discovery.
-We introduced a new benchmark to measure chemical exploration highlighting the poor intrinsic exploration of current baseline algorithms.
-We showed that this benchmark could be practically solved by a population size of up to 128 independent agents which performed superior to a single agent trained for longer, even with exploration bonuses.
-Lastly, the cooperative strategies we proposed and tested were unable to significantly outperform independent agents in our setting.
-The scaling of populations of agents in our setting has a direct relation to the scaling of TTT in LLMs.
-In our case, a CLM agent is pre-trained to form a generalist model that can sample molecules in a similar chemical space to the training dataset.
-We temporarily update the CLM agent with RL on a specific test-time task (each of the four benchmark tasks), which can be viewed as iterations of training and sampling, and collect the outcomes (molecules sampled) but throw away the updated parameters.
-However, as opposed to scaling just the inference-time as typically done, we scaled the whole TTT process by increasing the number of agents undergoing TTT on a specific test-time task.
-Interestingly, observing the same log-linear scaling law in performance with the number of samples.
-Overall, this work takes an important step towards the future scenario of identifying all possible molecules of interest in drug-like chemical space.
+In this paper, we probe the exploration of chemical space with a CLM to identify multiple candidate molecules in the context of drug discovery.
+In this effort, we introduce MolExp as a new benchmark, showing poor intrinsic exploration of current baseline algorithms.
+We showed that scaling a population of independent RL agents was the best approach to increasing benchmark performance, practically solving it with 128 agents.
+This included comparisons to scaling budgets for single RL agents and an extensive range of cooperative strategies.
+Moreover, we found that performance improved log-linearly, similarly observed with test-time inference scaling of LLMs.
+In this regard, training the pre-trained CLM agent with RL on a specific test-time task (each of the four benchmark tasks) can be viewed as TTT.
+We than scale the whole TTT process by increasing the number of RL agents, resulting in similar scaling laws observed with test-time inference scaling in LLMs.
+Overall, this work takes an important step towards identifying all possible molecules of interest in drug-like chemical space.
 Our MolExp benchmark is the first to explicitly ask this question, which we hope will guide algorithmic improvement until chemical space can be efficiently and thoroughly explored for a given context.
-Populations of RL agents are a promising solution towards this future.
-Impact statement
-This paper presents work whose goal is to advance the field of Machine Learning in application to drug discovery.
-We don’t anticipate any direct societal consequences of our work that must be specifically highlighted here.
-Appendix A Chemical language model implementation
-We trained a gated recurrent unit (GRU) network on molecules within the ChEMBL34 database (Gaulton et al., 2012) to act as the CLM.
-First, we filtered the database to form the training dataset and ensured no molecules existed with more 10 rotatable bonds, a logP above 5.5, a molecular weight outside of 150-650 Da, or contained atoms not in a comment set $a\in{C,N,S,O,F,Cl,Br,H}$.
+TTT scaling through populations of RL agents are a promising solution towards this future.
+Limitations
+The similarity functions used for the tasks are unlikely to be used in real-world drug design.
+In practice, predictive models of bioactivity, binding affinity, and ADMET properties are more common.
+However, the set of all the best possible solutions for such scoring functions (e.g., molecular docking) are currently unknowable.
+For this reason, MolExp(L) implements similarity functions as perfect scoring functions and tasks with known ground truths.
+We acknowledge this difference and strengthen the translational assumption with similar empirical results on other established benchmarks and a real-world objective in A2A bioactivity prediction.
+We did not investigate the effect of model architecture and training dataset in this work, which has been thoroughly investigated elsewhere.
+Instead we focus on maximizing chemical exploration efficiency given the same or similar starting policy.
+These results are expected to translate to any starting policy.
+The caveat of scaling TTT with more RL agents or increased budget is increased walltime requirements, as shown in Appendix I.
+This reiterates the importance of trying to use cooperative methods to improve exploration efficiency, as well as, the use of this benchmark and results as a tool for further research into targeted and efficient chemical space exploration.
+Code availability
+The code used to generate this data is part of the ACEGEN repository https://github.com/Acellera/acegen-open.
+The code and training dataset for the MolExp benchmark is part of the MolScore repository https://github.com/MorganCThomas/MolScore which is integrated into ACEGEN.
+Appendix A Model pre-training
+A.1 Datasets
+MolExp benchmark The pre-training dataset for the MolExp(L) benchmark was curated from the ChEMBL34 database [44].
+First, the database was filtered to ensure no molecules existed with more 10 rotatable bonds, a logP above 5.5, a molecular weight outside of 150-650 Da, or contained atoms not in a comment set $a\in{C,N,S,O,F,Cl,Br,H}$.
 All molecules were neutralized if possible, converted to non-isomeric SMILES (i.e., containing no stereoinformation) and de-duplicated.
@@
 Finally, the target molecules of the MolExp benchmark Appendix C were purposefully confirmed to be present within the training dataset, to ensure that objectives were theoretically achievable.
-The GRU consisted of an embedding of 256 and 3 layers of GRU cells with hidden dimension 512.
-This was trained with restricted SMILES randomization (Arús-Pous et al., 2019b) for 10 epochs with a batch size of 128.
-This was implemented such that a random SMILES string was each time a dataset molecule was sampled during training.
+GuacaMol benchmark The pre-training dataset for the GuacaMol [12] benchmark was taken directly from the original publication without any further processing.
+This dataset was downloaded from https://figshare.com/projects/GuacaMol/56639.
+A.2 ACEGEN
+MolExp benchmark The CLM constituting the prior policy is a gated recurrent unit (GRU) network with an embedding of 256 and 3 layers of GRU cells with hidden dimension 512.
+This was trained with restricted SMILES randomization [45] for 10 epochs with a batch size of 128.
+This was implemented such that a random SMILES string augmentation was applied each time a dataset molecule was sampled during training.
 An ADAM optimizer was employed with a learning rate of 0.001 and step scheduler that dropped the learning rate every 500 steps.
-The learning rate, loss and SMILES validity during training is shown in Figure 6.
-Figure 6: Pre-training of the CLM on a training dataset of ChEMBL34 molecules.
+The learning rate, training loss, and SMILES validity during training is shown in Figure 6.
+Figure 6: Pre-training of the GRU CLM on the MolExp benchmark pre-training dataset curated from ChEMBL34.
+GuacaMol benchmark The CLM constituting the prior policy is an long short-term memory (LSTM) network with the same hyperparameters as the model used in GuacaMol benchmark.
+This includes an embedding of 1024 and 3 layers of LSTM cells with hidden dimension 1024, as well as a dropout rate of 0.2.
+The pre-trained weights were loaded from the GuacaMol benchmark resulting in the exact same prior and initial policy.
+A.3 MolRL-MGPT
+The MolRL-MGPT model was re-implemented following the code provided from the original publication [36].
+This CLM constituting the prior policy is a GPT-style Transformer model with 8 layers each with 8 attention heads.
+For a standardized comparison, the model was pre-trained on the benchmark training datasets respectively.
+For hardware reasons the batch size was lowered from 2048 to 512, with all other hyperparameters kept equal including training for 10 epochs.
+The learning rate, training loss, and SMILES validity during training is shown in x.
+Figure 7: Pre-training of the MolRL-MGPT CLM on the MolExp benchmark pre-training dataset and GuacaMol benchmark pre-training dataset.
 Appendix B Cooperative strategies
@@
 As initially proposed by Hu et al.
-(2024), an additional loss penalty was added which seeks to maximize the difference in log-probability for on-policy states $\mathcal{S}^{k}$ between the $k$-th agent and previous agents, scaled by the reward $R(\tau)$.
+[36], an additional loss penalty was added which seeks to maximize the difference in log-probability for on-policy states $\mathcal{S}^{k}$ between the $k$-th agent and previous agents, scaled by the reward $R(\tau)$.
 $\mathcal{L}^{k}_{DIFF_{\mathcal{N}}}=-\lambda\cdot\sum_{i=1}^{k-1}R(\tau)\cdot\lvert\log\pi_{k}(\tau)-\log\pi_{i}(\tau)\rvert$ (12)
@@
 Taking inspiration from Parker-Holder et al.
-(2020) we implemented Diversity via Determinant to enforce divergent behavior of agent policies.
+[27], we implemented Diversity via Determinant to enforce divergent behavior of agent policies.
 A random selection of 100 on-policy states were sampled and used to form an embedding that encoded policy behavior.
@@
 $\mathcal{L}^{k}_{\text{POPNORM}}=-\frac{1}{k-1}\sum^{k-1}_{i=1}\pi_{i}(\tau)\cdot R(\tau)$ (15)
+Figure 8: Application of noise with varying coefficients $\lambda_{Noise}$ to either the final linear layer weights or all GRU hidden layers and the final linear layer weights.
+Increasing amounts of noise decrease the validity of molecules generated.
 Appendix C Molecular exploration benchmark
@@
 $R(m)=max(sim(m_{i},m_{t_{1}}),...,sim(m_{i},m_{t_{N}}))$ (16)
-In this work, the similarity function $sim$ is implemented as the Levenshtein similarity between the canonical SMILES strings.
-This choice is because the CLM generates and acts in a SMILES environment.
-We created the Levenshtein similarity by normalizing the Levenshtein distance by the length of the reference string, clipped to ensure no negative values:
+In this work, the similarity function $sim$ is implemented as the Levenshtein similarity between the canonical SMILES strings (MolExpL).
+This was chosen because the CLM generates and acts in a SMILES environment.
+The Levenshtein similarity function was created by normalizing the Levenshtein distance by the length of the reference string, clipped to ensure the output value was bound between [0, 1]:
 $sim_{\mathbb{L}}(m,m_{t})=clip(\frac{1-dis_{\mathbb{L}}(m,m_{t})}{|m_{t}|})$ (17)
 Note that the SMILES strings of both $m$ $m_{t}$ are canonicalized and both used non-isomeric SMILES.
-Moreover,we made an alternative similarity function available using the Tanimoto similarity of ECFP fingerprints for application with non-language models, although the Levenshtein similarity will still work provided a SMILES representation is given.
+Moreover,we made an alternative similarity function available using the Tanimoto similarity of ECFP fingerprints for application with non-language models (MolExp), although the Levenshtein similarity will still work provided a SMILES representation is given.
 Performance was evaluated based on the ability to generate molecules close to all molecular targets by taking the product:
@@
 All the molecular candidates are A2A receptor antagonists that bind to the same orthosteric site but with alternative binding modes.
-XAC binds from N253 up through the pocket with a polar tail that extends into solvent, LUAA47070 binds via a water mediated interaction with N253, while Imaradenant binds from N253 down occupying lipophilic hotspots
+XAC binds from N253 up towards the extracellular side with a polar tail that extends into solvent, LUAA47070 binds via a water mediated interaction with N253, while Imaradenant binds from N253 down towards the intrahelical bundle occupying lipophilic hotspots.
 Figure 11: Beta-secretase I (BACE1) receptor drug candidates as molecular targets.
@@
 $\nabla J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\cdot R(\tau)\right]$ (19)
-REINVENT: This baseline was implemented following the REINVENT formulation (Olivecrona et al., 2017) whereby a fixed version of the prior policy $\pi_{prior}$ (i.e., the pre-trained CLM) is used to regularize the updates of the agent policy $\pi_{agent}$ being trained.
+REINVENT: This baseline was implemented following the REINVENT formulation [14] whereby a fixed version of the prior policy $\pi_{prior}$ (i.e., the pre-trained CLM) is used to regularize the updates of the agent policy $\pi_{agent}$ being trained.
 This is a form of reward shaping as shown in following Equation 20.
@@
 All configuration parameter are shown in Table 4.
-[…]
+AHC: Augmented Hill-Climb [17] is a variation of REINVENT that conducts on-policy hill-climbing, i.e., on-policy molecules are ranked by reward $R(\tau)$ and only the top-$k$ molecules are retained.
+This shift focuses learning on high-reward molecules which improves efficiency, and helps to avoid too much regularization in low-reward scenarios.
+All configuration parameters are shown in Table 5.
+MolRL-MGPT: MolRL-MGPT is a cooperative multi-agent RL baseline with 4 GPT-style CLM agents.
+This follows the REINVENT formulation with the addition of a second loss term for the $k$-th agent encouraging reward scaled deviation of the current agent to previous ones.
+This baseline was implemented using the code provided alongside the publication [36].
+The default hyperparameters provided with the code were used, as there were insufficient details for full reproducibility of paper results.
+$\mathcal{L}^{k}_{DIFF_{\mathcal{N}}}=-\sigma_{2}\cdot\sum_{i=1}^{k-1}R(\tau)\cdot\lvert\log\pi_{k}(\tau)-\log\pi_{i}(\tau)\rvert$ (21)
 Table 2: Hyperparameters for REINFORCE.
@@
 baseline False
-Appendix E Computational resource
-All experiments detailed in this work were conducted on a single consumer grade GPU, more specifically an NVIDIA RTX 3090 with 1 CPU core.
-Figure 11: Wall time required when running the MolExp benchmark with increasing total budget sizes (larger population size for multi-agent RL).
-Appendix F Supplementary figures
-Figure 12: Application of noise with varying coefficients $\lambda_{Noise}$ to either the final linear layer weights or all GRU hidden layers and the final linear layer weights.
-Increasing amounts of noise decrease the validity of molecules generated.
+Table 8: Hyperparameters for MolRL-MGPT.
+Hyperparameter Value
+num_envs 128
+total_smiles 10,000
+model GPT (embedding of size 256 + 8 layer with 8 attention heads + MLP)
+lr 0.0001
+experience_replay True
+replay_batch_size 5
+replay_buffer_size 25
+replay_sampler uniform
+sigma1 100
+sigma2 0.5
+Appendix E Baseline performance
+E.1 MolExpL
+Figure 13: Baseline performance for each MolExpL task during RL training, single replicate.
+Each line color represents similarity to a target molecule.
+Note that ACEGEN and REINVENTMolOpt methods outperform due to their enhanced ability to optimize similarity to at-least one target molecule.
+Interestingly REINVENTMolOpt optimizes similarity to a different target molecule than ACEGEN for the EGFR task, despite identical initial policy parameterization.
+Figure 14: Estimated maximum performance achieved by ACEGENMolOpt.
+ACEGENMolOpt was trained to maximize the similarity to each individual molecular target separately for a budget of 10,000 molecules.
+Based on this, the estimated maximum performance on the benchmark is 2.92.
+E.2 MolExp
+Table 9: Baseline performance on the MolExp benchmark.
+SCREEN MolRL-MGPT REINFORCE REINVENT REINVENTMolOpt AHC ACEGENPractical ACEGENMolOpt
+AP 0.32 ± 0.10 0.24 ± 0.05 0.37 ± 0.02 0.25 ± 0.04 0.45 ± 0.03 0.47 ± 0.04 0.40 ± 0.02 0.42 ± 0.04
+A2A 0.08 ± 0.02 0.08 ± 0.01 0.09 ± 0.01 0.06 ± 0.01 0.12 ± 0.02 0.11 ± 0.03 0.13 ± 0.03 0.13 ± 0.05
+BACE1 0.16 ± 0.07 0.07 ± 0.01 0.05 ± 0.01 0.05 ± 0.01 0.06 ± 0.01 0.06 ± 0.01 0.06 ± 0.01 0.06 ± 0.01
+EGFR 0.04 ± 0.01 0.03 ± 0.01 0.04 ± 0.01 0.03 ± 0.00 0.05 ± 0.02 0.04 ± 0.01 0.05 ± 0.02 0.06 ± 0.01
+Sum 0.60 ± 0.12 0.42 ± 0.05 0.55 ± 0.02 0.39 ± 0.04 0.68 ± 0.05 0.68 ± 0.05 0.64 ± 0.04 1.07 ± 0.08
 Figure 15: Baseline performance for each MolExp task during RL training, single replicate.
-Each line represents similarity to a single objective or target molecule.
-Note that ACEGEN and REINVENTMolOpt methods outperform due to their enhanced ability to optimize similarity to at-least one objective.
-Interestingly REINVENTMolOpt optimizes similarity to a different target molecule than ACEGEN for the EGFR task, despite identical initial policy parameterization.
-Figure 14: Estimated maximum performance achieved by ACEGENMolOpt. ACEGENMolOpt was trained to maximize the similarity to each individual molecular target separately for a budget of 10,000 molecules.
-Figure 15: Multiple independent agent learning for each MolExp task during RL training, single replicate.
-Each line represents average similarity to a single objective or target molecule, and for clarity the standard deviation is not plotted due to already high variance across agents.
-As the number of agents increases (increasing the total budget), by natural stochasticity it is more likely that there is divergence in agent behavior.
-Interestingly, even with 128 independent agents all seem to choose optimization of one molecule of the other for the AP task.
-Figure 16: Single-agent learning for each MolExp task during RL training, single replicate.
-Each line represents similarity to a single objective or target molecule.
+Each line color represents similarity to a target molecule.
+Note that ACEGEN and REINVENTMolOpt methods outperform due to their enhanced ability to optimize similarity to at-least one target molecule.
+Figure 16: Estimated maximum performance achieved by ACEGENMolOpt.
+ACEGENMolOpt was trained to maximize the similarity to each individual molecular target separately for a budget of 10,000 molecules.
+Based on this, the estimated maximum performance on the benchmark is 2.41.
+E.3 GuacaMol
+Baseline algorithms were run on the GuacaMol benchmark also with 5 replicates, and for a standardized comparison, a budget constraint of 10,000 molecules.
+Previously published benchmark values have no restriction on budget, therefore, varying budgets have been used per algorithm.
+For this reason, values will differ from those previously published.
+Table 10: Baseline performance on the GuacaMol benchmark, part I.
+Task SMILES LSTM SMILES GA Graph GA Graph MCTS Frag GT SCREEN
+Albuterol similarity 0.92 ± 0.02 0.45 ± 0.05 0.91 ± 0.05 0.59 ± 0.03 0.99 ± 0.00 0.53 ± 0.01
+Amlodipine MPO 0.65 ± 0.02 0.52 ± 0.05 0.65 ± 0.04 0.44 ± 0.01 0.69 ± 0.04 0.56 ± 0.03
+Aripiprazole similarity 0.81 ± 0.03 0.37 ± 0.04 0.83 ± 0.10 0.25 ± 0.01 0.87 ± 0.06 0.50 ± 0.00
+C11H24 0.59 ± 0.05 0.01 ± 0.01 0.63 ± 0.16 0.06 ± 0.01 0.81 ± 0.02 0.05 ± 0.01
+C9H10N2O2PF2Cl 0.60 ± 0.01 0.14 ± 0.06 0.67 ± 0.10 0.56 ± 0.02 0.77 ± 0.03 0.39 ± 0.01
+Celecoxib rediscovery 1.00 ± 0.00 0.30 ± 0.01 0.70 ± 0.12 0.29 ± 0.03 0.70 ± 0.07 0.44 ± 0.03
+Deco hop 0.66 ± 0.01 0.59 ± 0.01 0.65 ± 0.02 0.55 ± 0.00 0.73 ± 0.01 0.62 ± 0.01
+Fexofenadine MPO 0.76 ± 0.01 0.67 ± 0.02 0.79 ± 0.01 0.58 ± 0.00 0.83 ± 0.02 0.72 ± 0.02
+Median molecules 1 0.34 ± 0.00 0.14 ± 0.02 0.28 ± 0.01 0.18 ± 0.01 0.28 ± 0.02 0.20 ± 0.01
+Median molecules 2 0.30 ± 0.01 0.18 ± 0.01 0.28 ± 0.02 0.13 ± 0.00 0.33 ± 0.06 0.23 ± 0.01
+Mestranol similarity 0.68 ± 0.05 0.32 ± 0.04 0.64 ± 0.11 0.27 ± 0.01 0.57 ± 0.02 0.48 ± 0.03
+Osimertinib MPO 0.83 ± 0.01 0.77 ± 0.02 0.83 ± 0.01 0.71 ± 0.01 0.87 ± 0.01 0.79 ± 0.00
+Perindopril MPO 0.51 ± 0.01 0.41 ± 0.03 0.53 ± 0.03 0.28 ± 0.01 0.58 ± 0.02 0.47 ± 0.02
+Ranolazine MPO 0.77 ± 0.01 0.55 ± 0.02 0.72 ± 0.02 0.24 ± 0.04 0.91 ± 0.01 0.71 ± 0.01
+Scaffold hop 0.58 ± 0.01 0.46 ± 0.01 0.58 ± 0.02 0.42 ± 0.00 0.77 ± 0.12 0.51 ± 0.02
+Sitagliptin MPO 0.34 ± 0.03 0.22 ± 0.03 0.49 ± 0.08 0.21 ± 0.02 0.45 ± 0.03 0.30 ± 0.02
+Thiothixene rediscovery 0.64 ± 0.04 0.30 ± 0.02 0.59 ± 0.07 0.25 ± 0.01 0.58 ± 0.05 0.41 ± 0.02
+Troglitazone rediscovery 0.48 ± 0.03 0.24 ± 0.02 0.44 ± 0.05 0.24 ± 0.01 0.45 ± 0.11 0.36 ± 0.03
+Valsartan smarts 0.03 ± 0.00 0.02 ± 0.00 0.03 ± 0.00 0.01 ± 0.00 0.03 ± 0.00 0.02 ± 0.00
+Zaleplon MPO 0.49 ± 0.01 0.36 ± 0.03 0.47 ± 0.01 0.32 ± 0.02 0.53 ± 0.02 0.45 ± 0.01
+Total Score 11.96 ± 0.10 7.01 ± 0.13 11.70 ± 0.30 6.59 ± 0.07 12.75 ± 0.22 8.76 ± 0.08
+Total Quality 16.03 ± 0.25 5.84 ± 0.56 9.96 ± 0.82 3.29 ± 0.29 13.40 ± 0.99 15.12 ± 0.19
+Table 11: Baseline performance on the GuacaMol benchmark, part II.
+Task MolRL-MGPT REINFORCE REINVENT REINVENTMolOpt AHC ACEGENPractical ACEGENMolOpt
+Albuterol similarity 0.60 ± 0.02 0.64 ± 0.03 0.60 ± 0.02 0.60 ± 0.02 0.71 ± 0.04 0.82 ± 0.05 1.00 ± 0.00
+Amlodipine MPO 0.58 ± 0.02 0.61 ± 0.01 0.59 ± 0.01 0.59 ± 0.01 0.63 ± 0.01 0.67 ± 0.02 0.84 ± 0.06
+Aripiprazole similarity 0.59 ± 0.06 0.58 ± 0.04 0.53 ± 0.02 0.53 ± 0.02 0.64 ± 0.03 0.74 ± 0.07 1.00 ± 0.00
+C11H24 0.04 ± 0.01 0.58 ± 0.04 0.06 ± 0.02 0.06 ± 0.02 0.06 ± 0.01 0.18 ± 0.08 0.88 ± 0.05
+C9H10N2O2PF2Cl 0.38 ± 0.02 0.54 ± 0.01 0.48 ± 0.01 0.48 ± 0.01 0.51 ± 0.01 0.48 ± 0.01 0.81 ± 0.02
+Celecoxib rediscovery 0.65 ± 0.13 0.56 ± 0.03 0.54 ± 0.07 0.54 ± 0.07 0.59 ± 0.03 0.87 ± 0.14 1.00 ± 0.00
+Deco hop 0.62 ± 0.01 0.62 ± 0.01 0.61 ± 0.01 0.61 ± 0.01 0.63 ± 0.00 0.64 ± 0.01 0.72 ± 0.03
+Fexofenadine MPO 0.73 ± 0.01 0.74 ± 0.02 0.74 ± 0.01 0.74 ± 0.01 0.73 ± 0.01 0.74 ± 0.01 0.87 ± 0.03
+Median molecules 1 0.20 ± 0.01 0.30 ± 0.02 0.25 ± 0.02 0.25 ± 0.02 0.28 ± 0.02 0.28 ± 0.04 0.39 ± 0.01
+Median molecules 2 0.21 ± 0.01 0.23 ± 0.01 0.23 ± 0.01 0.23 ± 0.01 0.25 ± 0.02 0.25 ± 0.02 0.36 ± 0.03
+Mestranol similarity 0.54 ± 0.05 0.58 ± 0.01 0.53 ± 0.01 0.53 ± 0.01 0.57 ± 0.02 0.76 ± 0.05 1.00 ± 0.00
+Osimertinib MPO 0.80 ± 0.01 0.81 ± 0.00 0.80 ± 0.00 0.80 ± 0.00 0.81 ± 0.00 0.80 ± 0.00 0.87 ± 0.01
+Perindopril MPO 0.48 ± 0.01 0.50 ± 0.00 0.48 ± 0.01 0.48 ± 0.01 0.49 ± 0.01 0.50 ± 0.01 0.60 ± 0.02
+Ranolazine MPO 0.75 ± 0.01 0.76 ± 0.01 0.75 ± 0.01 0.75 ± 0.01 0.76 ± 0.01 0.76 ± 0.00 0.85 ± 0.01
+Scaffold hop 0.52 ± 0.01 0.52 ± 0.01 0.52 ± 0.00 0.52 ± 0.00 0.53 ± 0.01 0.57 ± 0.03 0.78 ± 0.18
+Sitagliptin MPO 0.28 ± 0.02 0.33 ± 0.03 0.31 ± 0.02 0.31 ± 0.02 0.29 ± 0.02 0.29 ± 0.02 0.48 ± 0.03
+Thiothixene rediscovery 0.38 ± 0.02 0.46 ± 0.01 0.42 ± 0.01 0.42 ± 0.01 0.46 ± 0.03 0.61 ± 0.08 0.98 ± 0.06
+Troglitazone rediscovery 0.36 ± 0.03 0.39 ± 0.05 0.37 ± 0.03 0.37 ± 0.03 0.38 ± 0.03 0.44 ± 0.05 0.86 ± 0.20
+Valsartan smarts 0.02 ± 0.00 0.03 ± 0.00 0.02 ± 0.00 0.02 ± 0.00 0.02 ± 0.00 0.02 ± 0.00 0.03 ± 0.00
+Zaleplon MPO 0.45 ± 0.01 0.47 ± 0.01 0.47 ± 0.01 0.47 ± 0.01 0.47 ± 0.01 0.47 ± 0.01 0.55 ± 0.04
+Total Score 9.19 ± 0.16 10.23 ± 0.10 9.29 ± 0.09 9.29 ± 0.09 9.82 ± 0.09 10.88 ± 0.22 14.85 ± 0.30
+Total Quality 15.00 ± 0.17 15.69 ± 0.19 15.22 ± 0.16 15.22 ± 0.16 15.38 ± 0.20 15.67 ± 0.25 15.72 ± 0.85
+Appendix F TTT scaling
+F.1 MolExpL
+Table 12: Performance on MolExpL benchmark with increasing number of independent REINFORCE agents, each with a budget of 10,000.
+Task 1 2 4 8 16 32 64 128
+AP 0.56 ± 0.05 0.59 ± 0.03 0.67 ± 0.05 0.64 ± 0.02 0.67 ± 0.03 0.72 ± 0.02 0.72 ± 0.02 0.75 ± 0.03
+A2A 0.41 ± 0.06 0.42 ± 0.04 0.46 ± 0.05 0.50 ± 0.04 0.55 ± 0.04 0.54 ± 0.04 0.62 ± 0.05 0.63 ± 0.03
+BACE1 0.22 ± 0.02 0.25 ± 0.02 0.26 ± 0.02 0.29 ± 0.03 0.29 ± 0.02 0.32 ± 0.04 0.34 ± 0.02 0.37 ± 0.03
+EGFR 0.15 ± 0.02 0.17 ± 0.03 0.17 ± 0.02 0.21 ± 0.02 0.24 ± 0.01 0.24 ± 0.01 0.29 ± 0.02 0.32 ± 0.02
+Sum 1.34 ± 0.08 1.42 ± 0.06 1.57 ± 0.08 1.64 ± 0.05 1.75 ± 0.06 1.82 ± 0.06 1.97 ± 0.06 2.08 ± 0.06
+Table 13: Performance on MolExpL benchmark with increasing number of independent ACEGENMolOpt agents, each with a budget of 10,000.
+Task 1 2 4 8 16 32 64 128
+AP 0.62 ± 0.03 0.64 ± 0.02 0.70 ± 0.06 0.72 ± 0.03 0.70 ± 0.02 0.72 ± 0.01 0.75 ± 0.05 0.75 ± 0.02
+A2A 0.41 ± 0.07 0.51 ± 0.08 0.65 ± 0.08 0.70 ± 0.15 0.76 ± 0.10 0.87 ± 0.13 0.86 ± 0.11 0.99 ± 0.01
+BACE1 0.31 ± 0.07 0.40 ± 0.03 0.49 ± 0.11 0.53 ± 0.09 0.64 ± 0.07 0.77 ± 0.03 0.82 ± 0.02 0.90 ± 0.05
+EGFR 0.27 ± 0.06 0.34 ± 0.05 0.46 ± 0.11 0.57 ± 0.07 0.65 ± 0.10 0.77 ± 0.09 0.80 ± 0.05 0.89 ± 0.04
+Sum 1.62 ± 0.12 1.89 ± 0.10 2.31 ± 0.18 2.51 ± 0.19 2.74 ± 0.15 3.13 ± 0.16 3.24 ± 0.13 3.52 ± 0.07
+Table 14: Performance on MolExpL benchmark with increasing budget for a single ACEGENMolOpt agent.
+Task 10k 20k 40k 80k 160k 320k 640k 1280k
+AP 0.62 ± 0.03 0.64 ± 0.02 0.68 ± 0.05 0.64 ± 0.03 0.65 ± 0.03 0.65 ± 0.06 0.65 ± 0.04 0.65 ± 0.05
+A2A 0.41 ± 0.07 0.45 ± 0.06 0.45 ± 0.06 0.45 ± 0.06 0.45 ± 0.06 0.45 ± 0.06 0.46 ± 0.06 0.46 ± 0.06
+BACE1 0.31 ± 0.07 0.30 ± 0.07 0.35 ± 0.03 0.32 ± 0.02 0.32 ± 0.06 0.32 ± 0.07 0.33 ± 0.03 0.34 ± 0.06
+EGFR 0.27 ± 0.06 0.25 ± 0.04 0.28 ± 0.05 0.27 ± 0.03 0.26 ± 0.03 0.25 ± 0.02 0.24 ± 0.04 0.25 ± 0.06
+Sum 1.62 ± 0.12 1.64 ± 0.11 1.76 ± 0.10 1.68 ± 0.08 1.68 ± 0.09 1.67 ± 0.11 1.67 ± 0.09 1.70 ± 0.12
+Table 15: Performance on MolExpL benchmark with increasing budget for a single ACEGENMolOpt agent with a RND exploration bonus.
+Task 10k 20k 40k 80k 160k 320k 640k 1280k
+AP 0.65 ± 0.03 0.66 ± 0.03 0.65 ± 0.04 0.65 ± 0.01 0.65 ± 0.02 0.66 ± 0.04 0.65 ± 0.01 0.67 ± 0.03
+A2A 0.43 ± 0.03 0.45 ± 0.03 0.47 ± 0.06 0.47 ± 0.06 0.47 ± 0.06 0.47 ± 0.06 0.47 ± 0.06 0.47 ± 0.06
+BACE1 0.23 ± 0.01 0.38 ± 0.05 0.35 ± 0.06 0.32 ± 0.03 0.34 ± 0.03 0.37 ± 0.04 0.38 ± 0.05 0.37 ± 0.02
+EGFR 0.18 ± 0.02 0.26 ± 0.04 0.32 ± 0.04 0.32 ± 0.07 0.31 ± 0.04 0.30 ± 0.03 0.30 ± 0.03 0.29 ± 0.06
+Sum 1.50 ± 0.05 1.75 ± 0.08 1.79 ± 0.10 1.76 ± 0.09 1.76 ± 0.08 1.80 ± 0.09 1.81 ± 0.09 1.80 ± 0.09
+Table 16: Performance on MolExpL benchmark with increasing budget for a single ACEGENMolOpt agent with DF penalization.
+Task 10k 20k 40k 80k 160k 320k 640k 1280k
+AP 0.66 ± 0.02 0.66 ± 0.04 0.65 ± 0.03 0.64 ± 0.02 0.65 ± 0.05 0.64 ± 0.02 0.65 ± 0.05 0.65 ± 0.02
+A2A 0.50 ± 0.04 0.45 ± 0.04 0.50 ± 0.04 0.50 ± 0.04 0.50 ± 0.04 0.50 ± 0.04 0.50 ± 0.04 0.49 ± 0.02
+BACE1 0.29 ± 0.05 0.38 ± 0.08 0.34 ± 0.09 0.35 ± 0.04 0.33 ± 0.03 0.32 ± 0.08 0.40 ± 0.08 0.37 ± 0.06
+EGFR 0.23 ± 0.06 0.27 ± 0.03 0.27 ± 0.06 0.24 ± 0.03 0.25 ± 0.04 0.32 ± 0.04 0.29 ± 0.07 0.28 ± 0.05
+Sum 1.67 ± 0.09 1.77 ± 0.10 1.76 ± 0.12 1.72 ± 0.06 1.72 ± 0.08 1.78 ± 0.10 1.84 ± 0.12 1.79 ± 0.08
+Figure 17: Multiple independent ACEGENMolOpt agents on the MolExpL benchmark tasks, single replicate.
+Each line represents average similarity to one of the tasks target molecules, and for clarity the standard deviation is not plotted due to already high variance across agents.
+As the number of agents increases (increasing the total budget) stochasticity increases, resulting in divergence in agent behavior.
+Interestingly, even with 128 independent agents all seem to choose maximization towards one AP target molecule over the other.
+Figure 18: Single ACEGENMolOpt agent on the MolExpL benchmark tasks, single replicate.
+Each line represents average similarity to one of the tasks target molecules.
 In general, a single agent optimizes for similarity to one target molecule and doesn’t change throughout training.
-One notably exception exists for EGFR with a total budget of 640,000, where the agent seems to switch molecular target (in this case from green to blue).
-Figure 17: Single-agent with RND exploration bonus for each MolExp task during RL training, single replicate.
-Each line represents similarity to a single objective or target molecule.
+One notable exception exists for EGFR with a budget of 640,000, where the agent seems to switch molecular target (in this case from green to blue).
+Figure 19: Single ACEGENMolOpt agent with RND exploration bonus on the MolExpL benchmark tasks, single replicate.
+Each line represents average similarity to one of the tasks target molecules.
 In general, a single agent optimizes for similarity to one target molecule and doesn’t change throughout training, even with exploration bonuses.
-Note that drops in average score occur from a budget of 320,000 onward, but never this never changes the target molecule maximized.
-Figure 18: Cooperative-agent learning for each MolExp task during RL training, single replicate, part I.
-Each line represents similarity to a single objective or target molecule.
+Note that average score can drop from a budget of 320,000 onward, but doesn’t resulted in a switching of target molecule.
+Figure 20: Single ACEGENMolOpt agent with DF penalization on the MolExpL benchmark tasks, single replicate. Each line represents average similarity to one of the tasks target molecules.
+F.2 MolExp
+Table 17: Performance on MolExp benchmark with increasing number of independent ACEGENMolOpt agents, each with a budget of 10,000.
+Task 1 2 4 8 16 32 64 128
+AP 0.42 ± 0.04 0.43 ± 0.03 0.50 ± 0.04 0.50 ± 0.02 0.48 ± 0.03 0.51 ± 0.02 0.54 ± 0.01 0.60 ± 0.09
+A2A 0.13 ± 0.05 0.13 ± 0.02 0.15 ± 0.03 0.16 ± 0.02 0.28 ± 0.15 0.30 ± 0.12 0.44 ± 0.12 0.61 ± 0.23
+BACE1 0.06 ± 0.01 0.07 ± 0.01 0.09 ± 0.01 0.10 ± 0.02 0.11 ± 0.02 0.12 ± 0.02 0.15 ± 0.04 0.18 ± 0.04
+EGFR 0.06 ± 0.01 0.08 ± 0.04 0.12 ± 0.03 0.18 ± 0.02 0.21 ± 0.06 0.24 ± 0.04 0.31 ± 0.05 0.36 ± 0.03
+Sum 0.66 ± 0.07 0.71 ± 0.05 0.86 ± 0.06 0.94 ± 0.04 1.08 ± 0.16 1.17 ± 0.13 1.44 ± 0.14 1.75 ± 0.25
+(a) Scaling RL agents
+(b) Molecular diversity
+Figure 21: Performance on the MolExp benchmark with scaling.
+(a) Scaling the number of independent ACEGENMolOpt, each with a budget of 10,000.
+(b) The diversity of sampled compounds as measured by sphere exclusion diversity.
+Figure 22: Multiple independent ACEGENMolOpt agents on the MolExp benchmark tasks, single replicate.
+Each line represents average similarity to one of the tasks target molecules, and for clarity the standard deviation is not plotted due to already high variance across agents.
+F.3 GuacaMol
+Table 18: Performance on GuacaMol benchmark with increasing number of independent ACEGENMolOpt agents, each with a budget of 10,000.
+Task 1 2 4 8 16 32
+Albuterol similarity 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00
+Amlodipine MPO 0.84 ± 0.06 0.86 ± 0.03 0.89 ± 0.00 0.89 ± 0.00 0.90 ± 0.00 0.90 ± 0.00
+Aripiprazole similarity 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00
+C11H24 0.88 ± 0.05 0.92 ± 0.04 0.97 ± 0.02 0.99 ± 0.00 0.99 ± 0.00 1.00 ± 0.00
+C9H10N2O2PF2Cl 0.81 ± 0.02 0.83 ± 0.03 0.85 ± 0.01 0.88 ± 0.00 0.88 ± 0.00 0.89 ± 0.01
+Celecoxib rediscovery 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00
+Deco hop 0.72 ± 0.03 0.73 ± 0.02 0.73 ± 0.02 0.73 ± 0.01 0.75 ± 0.00 0.75 ± 0.00
+Fexofenadine MPO 0.87 ± 0.03 0.90 ± 0.02 0.89 ± 0.01 0.91 ± 0.01 0.92 ± 0.01 0.93 ± 0.01
+Median molecules 1 0.39 ± 0.01 0.41 ± 0.03 0.42 ± 0.01 0.44 ± 0.01 0.44 ± 0.01 0.44 ± 0.00
+Median molecules 2 0.36 ± 0.03 0.37 ± 0.02 0.39 ± 0.02 0.40 ± 0.01 0.41 ± 0.01 0.40 ± 0.00
+Mestranol similarity 1.00 ± 0.00 1.00 ± 0.01 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00
+Osimertinib MPO 0.87 ± 0.01 0.88 ± 0.01 0.88 ± 0.01 0.89 ± 0.01 0.89 ± 0.00 0.90 ± 0.01
+Perindopril MPO 0.60 ± 0.02 0.65 ± 0.03 0.65 ± 0.05 0.71 ± 0.05 0.72 ± 0.04 0.76 ± 0.01
+Ranolazine MPO 0.85 ± 0.01 0.85 ± 0.01 0.85 ± 0.00 0.85 ± 0.01 0.87 ± 0.01 0.87 ± 0.01
+Scaffold hop 0.78 ± 0.18 0.86 ± 0.13 0.79 ± 0.18 0.94 ± 0.06 0.99 ± 0.00 0.99 ± 0.01
+Sitagliptin MPO 0.48 ± 0.03 0.47 ± 0.04 0.52 ± 0.07 0.56 ± 0.04 0.58 ± 0.03 0.61 ± 0.06
+Thiothixene rediscovery 0.98 ± 0.06 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00
+Troglitazone rediscovery 0.86 ± 0.20 0.87 ± 0.17 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00
+Valsartan smarts 0.03 ± 0.00 0.03 ± 0.00 0.03 ± 0.00 0.03 ± 0.00 0.03 ± 0.00 0.03 ± 0.00
+Zaleplon MPO 0.55 ± 0.04 0.55 ± 0.00 0.58 ± 0.01 0.58 ± 0.02 0.61 ± 0.03 0.62 ± 0.03
+GuacaMol Score 14.85 ± 0.30 15.17 ± 0.23 15.47 ± 0.20 15.79 ± 0.10 15.98 ± 0.06 16.10 ± 0.07
+GuacaMol Quality 15.72 ± 0.85 15.05 ± 0.52 15.27 ± 0.53 14.53 ± 0.50 14.25 ± 0.44 13.99 ± 0.45
+(a) Scaling RL agents
+(b) Molecular diversity
+Figure 23: Performance on the GuacaMol benchmark with scaling.
+(a) Scaling the number of independent ACEGENMolOpt, each with a budget of 10,000.
+(b) The diversity of sampled compounds as measured by sphere exclusion diversity.
+Appendix G Cooperative RL
+G.1 MolExpL
+Table 19: Performance on MolExpL benchmark of different 4-agent cooperative strategies, part I.
+Task Independent Noise Shared Shared w. bonus Purge RND DvD POPNORM MolRL-MGPT
+AP 0.70 ± 0.06 0.66 ± 0.02 0.68 ± 0.04 0.64 ± 0.01 0.65 ± 0.02 0.68 ± 0.02 0.69 ± 0.04 0.67 ± 0.03 0.54 ± 0.02
+A2A 0.65 ± 0.08 0.64 ± 0.18 0.51 ± 0.12 0.55 ± 0.11 0.63 ± 0.05 0.66 ± 0.07 0.68 ± 0.10 0.68 ± 0.08 0.52 ± 0.05
+BACE1 0.49 ± 0.11 0.53 ± 0.10 0.39 ± 0.03 0.37 ± 0.03 0.42 ± 0.13 0.35 ± 0.06 0.40 ± 0.05 0.53 ± 0.08 0.31 ± 0.03
+EGFR 0.46 ± 0.11 0.44 ± 0.06 0.29 ± 0.06 0.29 ± 0.03 0.45 ± 0.06 0.41 ± 0.06 0.43 ± 0.05 0.48 ± 0.05 0.23 ± 0.04
+Sum 2.31 ± 0.18 2.27 ± 0.21 1.86 ± 0.14 1.86 ± 0.12 2.16 ± 0.15 2.10 ± 0.11 2.19 ± 0.13 2.35 ± 0.13 1.59 ± 0.07
+Table 20: Performance on MolExpL benchmark of different 4-agent cooperative strategies, part II.
+Task ENTS-0.001 ENTS-0.01 ENTS-0.1 ENTS-1 CES-0.001 CES-0.01 CES-0.1 CES-1
+AP 0.66 ± 0.03 0.67 ± 0.03 0.65 ± 0.02 0.68 ± 0.02 0.69 ± 0.03 0.68 ± 0.01 0.68 ± 0.04 0.68 ± 0.04
+A2A 0.67 ± 0.20 0.72 ± 0.12 0.65 ± 0.12 0.60 ± 0.08 0.65 ± 0.18 0.62 ± 0.07 0.73 ± 0.13 0.70 ± 0.09
+BACE1 0.59 ± 0.08 0.51 ± 0.17 0.50 ± 0.09 0.41 ± 0.05 0.46 ± 0.15 0.53 ± 0.08 0.43 ± 0.05 0.36 ± 0.07
+EGFR 0.47 ± 0.05 0.44 ± 0.10 0.47 ± 0.09 0.38 ± 0.07 0.46 ± 0.15 0.44 ± 0.10 0.47 ± 0.08 0.39 ± 0.11
+Sum 2.39 ± 0.22 2.33 ± 0.23 2.28 ± 0.18 2.07 ± 0.12 2.25 ± 0.28 2.27 ± 0.14 2.31 ± 0.16 2.13 ± 0.16
+Table 21: Performance on MolExpL benchmark of different 4-agent cooperative strategies, part III.
+Task DIFFS-0.001 DIFFS-0.01 DIFFS-0.1 DIFFS-1 DIFFN-0.001 DIFFN-0.01 DIFFN-0.1 DIFFN-1
+AP 0.68 ± 0.02 0.69 ± 0.02 0.66 ± 0.03 0.63 ± 0.03 0.67 ± 0.03 0.68 ± 0.03 0.66 ± 0.02 0.68 ± 0.04
+A2A 0.57 ± 0.07 0.61 ± 0.08 0.54 ± 0.06 0.46 ± 0.04 0.68 ± 0.14 0.64 ± 0.04 0.63 ± 0.12 0.74 ± 0.07
+BACE1 0.45 ± 0.06 0.46 ± 0.06 0.35 ± 0.04 0.31 ± 0.06 0.48 ± 0.11 0.43 ± 0.13 0.47 ± 0.07 0.47 ± 0.13
+EGFR 0.45 ± 0.10 0.44 ± 0.05 0.36 ± 0.10 0.26 ± 0.05 0.50 ± 0.04 0.53 ± 0.03 0.45 ± 0.05 0.50 ± 0.10
+Sum 2.14 ± 0.13 2.20 ± 0.12 1.92 ± 0.13 1.66 ± 0.10 2.33 ± 0.19 2.27 ± 0.14 2.21 ± 0.15 2.39 ± 0.18
+Figure 24: Cooperative strategies using 4-agents on the MolExpL benchmark tasks, single replicate, part I.
+Each line represents average similarity to one of the tasks target molecules.
 Training of each agent is plotted.
 Perfect cooperation is each agent learning to rediscover a different target molecule.
-Most cooperative strategies investigated result in no additional divergent behavior or in some cases, slower learning.
-Figure 19: Cooperative-agent learning for each MolExp task during RL training, single replicate, part II.
-Each line represents similarity to a single objective or target molecule.
+Most cooperative strategies result in no additional divergent behavior or in some cases, slower learning.
+Figure 25: Cooperative strategies using 4-agents on the MolExpL benchmark tasks, single replicate, part II.
+Each line represents average similarity to one of the tasks target molecules.
 Training of each agent is plotted.
 Perfect cooperation is each agent learning to rediscover a different target molecule.
-Most cooperative strategies investigated result in no additional divergent behavior or in some cases, slower learning.
-Figure 20: Cooperative-agent learning for each MolExp task during RL training, single replicate, part III.
-Each line represents similarity to a single objective or target molecule.
+Most cooperative strategies result in no additional divergent behavior or in some cases, slower learning.
+Figure 26: Cooperative strategies using 4-agents on the MolExpL benchmark tasks, single replicate, part III.
+Each line represents average similarity to one of the tasks target molecules.
 Training of each agent is plotted.
 Perfect cooperation is each agent learning to rediscover a different target molecule.
-Most cooperative strategies investigated result in no additional divergent behavior or in some cases, slower learning.
+Most cooperative strategies result in no additional divergent behavior or in some cases, slower learning.
+G.2 MolExp
+Table 22: Performance on MolExp benchmark of different 4-agent cooperative strategies, part I.
+Task Independent Noise Shared Shared w. bonus Purge RND DvD POPNORM MolRL-MGPT
+AP 0.50 ± 0.04 0.47 ± 0.04 0.43 ± 0.02 0.44 ± 0.02 0.44 ± 0.01 0.47 ± 0.04 0.48 ± 0.06 0.46 ± 0.03 0.31 ± 0.06
+A2A 0.15 ± 0.03 0.25 ± 0.13 0.15 ± 0.04 0.17 ± 0.03 0.17 ± 0.05 0.17 ± 0.04 0.20 ± 0.12 0.17 ± 0.06 0.12 ± 0.02
+BACE1 0.09 ± 0.01 0.09 ± 0.01 0.07 ± 0.01 0.07 ± 0.01 0.08 ± 0.01 0.08 ± 0.01 0.07 ± 0.00 0.08 ± 0.01 0.15 ± 0.04
+EGFR 0.12 ± 0.03 0.11 ± 0.04 0.07 ± 0.02 0.06 ± 0.01 0.12 ± 0.04 0.08 ± 0.01 0.11 ± 0.02 0.13 ± 0.10 0.05 ± 0.01
+Sum 1.30 ± 0.07 0.91 ± 0.14 0.73 ± 0.05 0.74 ± 0.04 0.81 ± 0.07 0.80 ± 0.06 0.86 ± 0.13 0.85 ± 0.12 0.62 ± 0.07
+Table 23: Performance on MolExp benchmark of different 4-agent cooperative strategies, part II.
+Task ENTS-0.001 ENTS-0.01 ENTS-0.1 ENTS-1 CES-0.001 CES-0.01 CES-0.1 CES-1
+AP 0.45 ± 0.04 0.45 ± 0.03 0.54 ± 0.07 0.47 ± 0.04 0.51 ± 0.04 0.49 ± 0.06 0.54 ± 0.09 0.48 ± 0.05
+A2A 0.15 ± 0.02 0.25 ± 0.12 0.23 ± 0.16 0.25 ± 0.13 0.22 ± 0.11 0.15 ± 0.03 0.42 ± 0.07 0.27 ± 0.12
+BACE1 0.11 ± 0.06 0.08 ± 0.00 0.09 ± 0.02 0.09 ± 0.01 0.08 ± 0.02 0.08 ± 0.01 0.11 ± 0.03 0.09 ± 0.01
+EGFR 0.13 ± 0.05 0.13 ± 0.04 0.12 ± 0.05 0.10 ± 0.03 0.10 ± 0.03 0.12 ± 0.04 0.14 ± 0.04 0.08 ± 0.04
+Sum 0.83 ± 0.08 0.91 ± 0.12 0.98 ± 0.18 0.91 ± 0.14 0.91 ± 0.13 0.84 ± 0.08 1.21 ± 0.12 0.92 ± 0.14
+Table 24: Performance on MolExp benchmark of different 4-agent cooperative strategies, part III.
+Task DIFFS-0.001 DIFFS-0.01 DIFFS-0.1 DIFFS-1 DIFFN-0.001 DIFFN-0.01 DIFFN-0.1 DIFFN-1
+AP 0.47 ± 0.04 0.45 ± 0.05 0.45 ± 0.03 0.42 ± 0.03 0.46 ± 0.03 0.47 ± 0.02 0.45 ± 0.01 0.53 ± 0.12
+A2A 0.16 ± 0.04 0.17 ± 0.04 0.14 ± 0.01 0.14 ± 0.02 0.16 ± 0.03 0.17 ± 0.04 0.19 ± 0.10 0.41 ± 0.19
+BACE1 0.09 ± 0.01 0.11 ± 0.03 0.07 ± 0.02 0.06 ± 0.01 0.07 ± 0.01 0.09 ± 0.02 0.11 ± 0.03 0.10 ± 0.03
+EGFR 0.12 ± 0.02 0.12 ± 0.01 0.09 ± 0.03 0.06 ± 0.02 0.10 ± 0.05 0.16 ± 0.03 0.11 ± 0.04 0.12 ± 0.02
+Sum 0.84 ± 0.06 0.84 ± 0.07 0.74 ± 0.05 0.69 ± 0.04 0.79 ± 0.06 0.89 ± 0.06 0.86 ± 0.12 1.15 ± 0.23
+(a) Cooperative RL MolExp benchmark performance
+(b) Cooperative RL molecular diversity
+Figure 27: Performance comparison of different 4-agent cooperative strategies on the MolExp benchmark, each with a budget of 10k.
+The dashed line represents the average of 4 independent agents as baseline.
+Figure 28: Cooperative strategies using 4-agents on the MolExp benchmark tasks, single replicate, part I.
+Each line represents average similarity to one of the tasks target molecules.
+Training of each agent is plotted.
+Perfect cooperation is each agent learning to rediscover a different target molecule.
+Most cooperative strategies result in no additional divergent behavior or in some cases, slower learning.
+Figure 29: Cooperative strategies using 4-agents on the MolExp benchmark tasks, single replicate, part II.
+Each line represents average similarity to one of the tasks target molecules.
+Training of each agent is plotted.
+Perfect cooperation is each agent learning to rediscover a different target molecule.
+Most cooperative strategies result in no additional divergent behavior or in some cases, slower learning.
+Figure 30: Cooperative strategies using 4-agents on the MolExp benchmark tasks, single replicate, part III.
+Each line represents average similarity to one of the tasks target molecules.
+Training of each agent is plotted.
+Perfect cooperation is each agent learning to rediscover a different target molecule.
+Most cooperative strategies result in no additional divergent behavior or in some cases, slower learning.
+Appendix H A2A bioactivity maximization
+Table 25: Performance on MolExpBio task with increasing number of independent ACEGENMolOpt agents, each with a budget of 10,000.
+1 2 4 8 16 32 64 128
+Target 1 max sim 0.31 ± 0.02 0.32 ± 0.03 0.37 ± 0.03 0.36 ± 0.01 0.43 ± 0.07 0.44 ± 0.04 0.49 ± 0.07 0.54 ± 0.07
+Target 2 max sim 0.31 ± 0.03 0.34 ± 0.03 0.36 ± 0.01 0.38 ± 0.03 0.39 ± 0.02 0.41 ± 0.02 0.43 ± 0.05 0.44 ± 0.04
+Target 3 max sim 0.50 ± 0.08 0.53 ± 0.06 0.55 ± 0.02 0.60 ± 0.03 0.60 ± 0.03 0.62 ± 0.03 0.63 ± 0.03 0.63 ± 0.01
+MolExpBio score 0.05 ± 0.00 0.06 ± 0.01 0.07 ± 0.01 0.08 ± 0.01 0.10 ± 0.02 0.11 ± 0.01 0.13 ± 0.02 0.15 ± 0.02
+Table 26: Performance on MolExpBio task with increasing budget for a single ACEGENMolOpt agent.
+10k 20k 40k 80k 160k 320k 640k 1280k
+Target 1 max sim 0.31 ± 0.02 0.31 ± 0.02 0.31 ± 0.02 0.31 ± 0.02 0.31 ± 0.02 0.31 ± 0.02 0.31 ± 0.02 0.31 ± 0.02
+Target 2 max sim 0.31 ± 0.03 0.31 ± 0.03 0.31 ± 0.03 0.31 ± 0.03 0.31 ± 0.03 0.31 ± 0.03 0.31 ± 0.03 0.31 ± 0.03
+Target 3 max sim 0.50 ± 0.08 0.50 ± 0.08 0.50 ± 0.08 0.50 ± 0.08 0.50 ± 0.08 0.50 ± 0.08 0.50 ± 0.08 0.50 ± 0.08
+MolExpBio score 0.05 ± 0.00 0.05 ± 0.00 0.05 ± 0.00 0.05 ± 0.00 0.05 ± 0.00 0.05 ± 0.00 0.05 ± 0.00 0.05 ± 0.00
+Table 27: Performance on MolExpBio task with increasing budget for a single ACEGENMolOpt agent with RND.
+10k 20k 40k 80k 160k 320k 640k 1280k
+Target 1 max sim 0.37 ± 0.07 0.37 ± 0.07 0.37 ± 0.07 0.37 ± 0.07 0.37 ± 0.07 0.34 ± 0.02 0.37 ± 0.07 0.37 ± 0.07
+Target 2 max sim 0.33 ± 0.02 0.33 ± 0.02 0.33 ± 0.02 0.33 ± 0.02 0.33 ± 0.02 0.34 ± 0.03 0.33 ± 0.02 0.33 ± 0.02
+Target 3 max sim 0.55 ± 0.06 0.56 ± 0.06 0.56 ± 0.06 0.56 ± 0.06 0.56 ± 0.06 0.54 ± 0.04 0.56 ± 0.06 0.56 ± 0.06
+MolExpBio score 0.07 ± 0.02 0.07 ± 0.02 0.07 ± 0.02 0.07 ± 0.02 0.07 ± 0.02 0.06 ± 0.01 0.07 ± 0.02 0.07 ± 0.02
+Table 28: Performance on MolExpBio task with increasing budget for a single ACEGENMolOpt agent with DF penalization.
+10k 20k 40k 80k 160k 320k 640k 1280k
+Target 1 max sim 0.31 ± 0.02 0.31 ± 0.02 0.31 ± 0.02 0.31 ± 0.02 0.31 ± 0.02 0.31 ± 0.02 0.31 ± 0.02 0.31 ± 0.02
+Target 2 max sim 0.31 ± 0.03 0.31 ± 0.03 0.31 ± 0.03 0.31 ± 0.03 0.31 ± 0.03 0.32 ± 0.02 0.32 ± 0.02 0.32 ± 0.02
+Target 3 max sim 0.50 ± 0.05 0.50 ± 0.05 0.50 ± 0.05 0.50 ± 0.05 0.50 ± 0.05 0.50 ± 0.05 0.50 ± 0.05 0.50 ± 0.05
+MolExpBio score 0.05 ± 0.00 0.05 ± 0.00 0.05 ± 0.00 0.05 ± 0.00 0.05 ± 0.00 0.05 ± 0.00 0.05 ± 0.00 0.05 ± 0.00
+Figure 31: Multiple independent ACEGENMolOpt agents on the MolExpBio task of maximizing A2A predicted bioactivity, single replicate.
+Maximization of predicted A2A bioactivity is shown, as well as the maximum similarity to each A2A target molecule in the set from the MolExp A2A task in red.
+Figure 32: Single ACEGENMolOpt agent on the MolExpBio task of maximizing A2A predicted bioactivity, single replicate.
+Maximization of predicted A2A bioactivity is shown, as well as the maximum similarity to each A2A target molecule in the set from the MolExp A2A task in red.
+Figure 33: Single ACEGENMolOpt agent with RND exploration bonus on the MolExpBio task of maximizing A2A predicted bioactivity, single replicate.
+Maximization of predicted A2A bioactivity is shown, as well as the maximum similarity to each A2A target molecule in the set from the MolExp A2A task in red.
+Figure 34: Single ACEGENMolOpt agent with DF penalization on the MolExpBio task of maximizing A2A predicted bioactivity, single replicate.
+Maximization of predicted A2A bioactivity is shown, as well as the maximum similarity to each A2A target molecule in the set from the MolExp A2A task in red.
+Appendix I Computational resource
+All experiments detailed in this work were conducted on a single consumer grade GPU, more specifically an NVIDIA RTX 3090 with 1 CPU core.
+(a) MolExpL benchmark
+(b) MolExp benchmark
+(c) GuacaMol benchmark
+Figure 35: Wall time required to run benchmarks detailed in this work. Not that the DF was implemented following some code optimizations.